### v2

# Affine Chabauty I

Abstract.
We prove finiteness and give an explicit upper bound on the number of $S$-integral points on affine curves satisfying a certain rank-genus inequality. We achieve this by developing an analogue of the Chabauty method, embedding the curve into its generalised Jacobian and bounding the Abel–Jacobi image of the $S$-integral points using arithmetic intersection theory. Our results also provide the foundations for a computational method to determine the set of $S$-integral points on affine curves which will be presented in a follow-up article.
1. Introduction
Let $Y/\mathbb{Q}$ be a smooth affine curve and let $\mathcal{Y}/\mathbb{Z}_{S}$ be a regular model of $Y$ over the ring of $S$-integers for some finite set of primes $S$. If $Y$ is hyperbolic, the set of $S$-integral points $\mathcal{Y}(\mathbb{Z}_{S})$ is finite by the theorems of Siegel, Mahler, and Faltings. However, the natural questions of finding the points in $\mathcal{Y}(\mathbb{Z}_{S})$ or bounding the size of this set remain difficult open problems in general. We present a new approach addressing these questions, which results in an explicit upper bound on $\mathcal{Y}(\mathbb{Z}_{S})$ for curves satisfying the hypothesis
$r+\#S<g+\#|D|+n_{2}(D)-1,$ (1.1)
an inequality involving the genus $g$, the Mordell–Weil rank $r$, the cardinality of $S$, the number of closed points and the number of conjugate pairs of complex points in the boundary $D$ of $Y$. Our method is an $S$-integral analogue of the method of Chabauty–Coleman [Cha41, Col85]. Here, we embed $Y$ into its generalised Jacobian $J_{Y}$, choose an auxiliary prime $p\not\in S$ (possibly of bad reduction), and cut out $\mathcal{Y}(\mathbb{Z}_{S})$ inside $\mathcal{Y}(\mathbb{Z}_{p})$ by $p$-adic integrals of logarithmic differentials, i.e. meromorphic differentials on a compactification of $Y$ with at worst simple poles at $D$. The existence of suitable log differentials is guaranteed by assumption (1.1), which ensures that the image of $\mathcal{Y}(\mathbb{Z}_{S})$ inside $J_{Y}(\mathbb{Q})$ is contained in a finite union of translates of subgroups of controllable rank. We then bound the size of $\mathcal{Y}(\mathbb{Z}_{S})$ by estimating the number of zeros of $p$-adic integrals of log differentials. In a follow-up article [LL], we present an algorithm to determine the log differentials and the zeros of their integrals in practice, thus computing a finite subset of $\mathcal{Y}(\mathbb{Z}_{p})$ containing $\mathcal{Y}(\mathbb{Z}_{S})$.
1.1. Main results
In this introduction, we work with curves over $\mathbb{Q}$, but in the body of the paper we formulate our results over number fields. Let $Y/\mathbb{Q}$ be given as $X\smallsetminus D$ where $X/\mathbb{Q}$ is a smooth projective curve and $D\neq\emptyset$ is a finite set of closed points called cusps. Let $S$ be a finite set of primes. Let $\mathcal{X}$ be a regular model of $X$ over the ring $\mathbb{Z}_{S}$ of $S$-integers, i.e., a regular, flat, projective $\mathbb{Z}_{S}$-scheme with an isomorphism $\mathcal{X}_{\mathbb{Q}}\cong X$ [Liu06, Definition 10.1.1]. Let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}\coloneqq\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. We use the following notation, which will be kept throughout this paper:
  • $r\coloneqq\operatorname{\mathrm{rk}}J(\mathbb{Q})$ the Mordell–Weil rank of the Jacobian $J$ of $X$;
  • $g$ the genus of $X$;
  • #|D|>0\#|D|>0 the number of cusps;
  • $n\coloneqq\#D({\overline{\mathbb{Q}}})>0$ the number of geometric cusps;
  • write $n=n_{1}(D)+2n_{2}(D)$ with $n_{1}(D)\coloneqq\#D(\mathbb{R})$ the number of real cusps and $n_{2}(D)$ the number of conjugate pairs of complex cusps.
We partition the $S$-integral points
$\mathcal{Y}(\mathbb{Z}_{S})=\coprod_{\Sigma}\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma},$
where $\Sigma$ runs through the finitely many $S$-integral reduction types of $\mathcal{Y}$, see ˜3.8. Roughly, an $S$-integral reduction type $\Sigma=(\Sigma_{\ell})_{\ell}$ specifies for each prime $\ell\not\in S$ a component $\Sigma_{\ell}$ of the mod-$\ell$ fibre $\mathcal{X}_{\ell}$, while for $\ell\in S$ it specifies either a component of the mod-$\ell$ fibre or the mod-$\ell$ reduction of a cusp on a suitable extension of $\mathcal{X}$ to a regular model over $\mathbb{Z}$. The subset $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}\subseteq\mathcal{Y}(\mathbb{Z}_{S})$ consists of those $S$-integral points that reduce onto $\Sigma_{\ell}$ modulo $\ell$ for all $\ell$. Our main result is:
Theorem A.
Let $\Sigma$ be an $S$-integral reduction type and let $C(\Sigma)\subseteq S$ denote the set of those $\ell\in S$ where $\Sigma_{\ell}$ is cuspidal, see ˜3.10. Let $p\not\in S$ be a prime. Assume
$r+\#C(\Sigma)<g+\#|D|+n_{2}(D)-1.$ (1.2)
Then there exist a non-zero log differential $0\neq\omega\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))$ and a constant $c\in\mathbb{Q}_{p}$ such that the function $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$ given by
$\rho(P)\coloneqq\int_{P_{0}}^{P}\omega-c$ (1.3)
vanishes on $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$.
˜A is a special case of ˜5.1, which holds over general number fields. As (1.1) implies (1.2) for every $S$-integral reduction type, we deduce:
Theorem B.
If the inequality (1.1) holds, then $\mathcal{Y}(\mathbb{Z}_{S})$ is contained in a finite union of zero loci of $p$-adic analytic functions on $\mathcal{Y}(\mathbb{Z}_{p})$.
The $p$-adic analytic functions (1.3) are integrals of log differentials plus a constant, and by analysing their Newton polygons we arrive at bounds for the number of $S$-integral points.
Theorem C.
If (1.1) holds and $p>2g+n$, then
$\mathcal{Y}(\mathbb{Z}_{S})\leq(\#\mathcal{Y}^{\mathrm{sm}}_{p}(\operatorname{\mathbb{F}}_{p})+2g-2+n)\prod_{\ell\in S}(n_{\ell}+\#(\mathcal{X}_{{\ell}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell}))\prod_{\ell\not\in S\cup\{p\}}n_{\ell},$
where $n_{\ell}$ denotes the number of components of $\mathcal{X}_{{\ell}}\smallsetminus\operatorname{\mathcal{D}}$ that contain a smooth $\operatorname{\mathbb{F}}_{\ell}$-point and $\mathcal{A}^{\mathrm{sm}}_{{\ell}}$ denotes the smooth locus of the special fibre of $\mathcal{A}\in\{\mathcal{X},\mathcal{Y}\}$ at $\ell$.
In fact, different reduction types might give rise to the same Chabauty function $\rho$ in (1.3), so the factor in ˜C coming from summing over all reduction types can be improved, see ˜6.7. We demonstrate this in ˜7.1 by exhibiting a large class of even degree hyperelliptic curves $Y\colon y^{2}=f(x)$ for which the rank condition $r=g$ implies the simple bound
$\mathcal{Y}(\mathbb{Z})\leq\#\mathcal{Y}(\operatorname{\mathbb{F}}_{p})+2g$
for any prime $p>2g+2$ of good reduction. We also include examples where the bound is sharp.
Remark 1.1.
The results in this article are logically independent from, but inspired by the following circle of ideas:
  1. (1)
    Chabauty’s method for projective curves [Cha41, Col85, MP12];
  2. (2)
    the “linear Chabauty” (or depth 1) part of Kim’s nonabelian Chabauty method for affine curves [Kim05, Kim09, BDCKW18];
  3. (3)
    the Chabauty–Skolem method [Sko34, Poo21, Tri21] for affine curves.
Our approach is closest to (1), but replaces the Jacobian by the generalised Jacobian, holomorphic differentials by logarithmic differentials, and uses arithmetic intersection theory to constrain the Abel–Jacobi image, see Section˜1.2.
Using approach (2), we previously showed in joint work with Steffen Müller that hypothesis (1.1) implies finiteness of the depth-1 Chabauty–Kim locus [LLM23, Theorem A(1)+Remark 1.4], and obtained bounds under a more restrictive hypothesis [LLM23, Remark 1.2]. The present manuscript improves upon this in several aspects: the bound in ˜C is smaller, we only need hypothesis (1.1) to prove both finiteness and a bound, we allow $p$ to be a prime of bad reduction for $\mathcal{X}$, our method works over general number fields, and we determine the type of Coleman functions on $\mathcal{Y}(\mathbb{Z}_{p})$ vanishing on $S$-integral points. Crucial for these improvements is the geometry and arithmetic of the generalised Jacobian. Moreover, the functions $\rho$ in ˜A can be algorithmically determined in practice, and by computing their zeros one finds a finite subset of $\mathcal{Y}(\mathbb{Z}_{p})$ containing $\mathcal{Y}(\mathbb{Z}_{S})$. In the case of an even-degree hyperelliptic curve over $\mathbb{Q}$ and $S=\emptyset$, such a method was recently developed in [GM25] using $p$-adic Coleman–Gross heights. This “Linear Quadratic Chabauty” method arises as a special case of our method, see Section 7.1.
Our method can also be seen as a refined version of (3) in the sense of [BD20]. Namely, the idea of (3) is to embed $\mathcal{Y}$ into a suitable model $\mathcal{J}/\mathbb{Z}_{S}$ of the generalised Jacobian $J_{Y}$. Finiteness of $\mathcal{Y}(\mathbb{Z}_{S})$ is ensured by
$\operatorname{\mathrm{rk}}\mathcal{J}(\mathbb{Z}_{S})<\dim J_{Y},$
which is usually stricter than our Affine Chabauty Condition (1.1), as can be seen by looking at the complement in $\mathbb{P}^{1}$ of $n\geq 2$ rational points, for example. The reason is ultimately that we get a more refined control on the Abel–Jacobi image in the generalised Jacobian by also incorporating local conditions at primes inside $S$ via the notion of $S$-integral reduction type.
Remark 1.2.
This article does not touch upon the “quadratic Chabauty” results in [LLM23]. There, the hypothesis for finiteness of the depth-2 locus $\mathcal{Y}(\mathbb{Z}_{p})_{S,2}^{\mathrm{BD}}$ has an additional term of $\rho_{f}=\operatorname{\mathrm{rk}}\operatorname{NS}(J)+\operatorname{\mathrm{rk}}\operatorname{NS}(J_{{\overline{\mathbb{Q}}}})^{\sigma=-1}$ on the right hand side of (1.1). We believe the methods of this article to be a first step in the direction of making affine quadratic Chabauty explicit as well, and aim to pursue this line of investigation in the future.
1.2. Strategy
Let $J_{Y}$ denote the generalised Jacobian of $(X,D)$, see [Ser88, Chapter V]. We have an Abel–Jacobi map $\operatorname{AJ}_{P_{0}}\colon Y\to J_{Y}$ associated to the base point $P_{0}$, which yields the Affine Chabauty diagram
$\mathcal{Y}(\mathbb{Z}_{S})$$\mathcal{Y}(\mathbb{Z}_{p})$$J_{Y}(\mathbb{Q})$$J_{Y}(\mathbb{Q}_{p})$$\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))^{\vee},$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\int_{P_{0}}}$$\scriptstyle{\log_{J_{Y}}}$ (1.4)
where $\Omega^{1}(D)$ is the module of logarithmic differentials on $(X,D)$. Unlike the situation for projective curves, where we have the Jacobian $J$ in the bottom row whose Mordell–Weil group $J(\mathbb{Q})$ is finitely generated, the generalised Jacobian $J_{Y}$ is only a semi-abelian variety and the group $J_{Y}(\mathbb{Q})$ is almost never finitely generated. The crucial observation is that if we only look at $S$-integral points of a given reduction type $\Sigma$, then the Abel–Jacobi image of $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ inside $J_{Y}(\mathbb{Q})$ does in fact lie in a translate of a certain finitely generated subgroup. The hypothesis (1.2) ensures that the rank of this subgroup is strictly less than $\dim_{\mathbb{Q}_{p}}\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))=g+n-1$, and from this point onwards we deduce ˜A by the “usual Chabauty logic”. Namely, the image of this subgroup under $\log_{J_{Y}}$ is contained in a proper $\mathbb{Q}_{p}$-subspace of $\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))^{\vee}$, so there is a non-trivial log differential $\omega$ vanishing on this image. The map $\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$, $P\mapsto\int_{P_{0}}^{P}\omega$ is then constant on $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$.
It remains to explain how we restrict the Abel–Jacobi image of $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ inside $J_{Y}(\mathbb{Q})$. We achieve this by constructing a homomorphism $\sigma\colon J_{Y}(\mathbb{Q})\to V_{D}$ to a certain $\mathbb{Q}$-vector space $V_{D}$ with the property that $\ker(\sigma)$ has finite rank and the image of $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ in $V_{D}$ is contained in a translate of a subgroup of small rank. The map $\sigma$ is constructed using arithmetic intersection theory on a regular model. For simplicity, assume that $\mathcal{X}$ is a regular model of $X$ over $\mathbb{Z}$ such that the closure $\operatorname{\mathcal{D}}$ of $D$ in $\mathcal{X}$ is normal. Consider the group $Z_{0}(\operatorname{\mathcal{D}})$ of zero-cycles on $\operatorname{\mathcal{D}}$ and let $\pi$ denote the map $\operatorname{\mathcal{D}}\to\operatorname{Spec}(\mathbb{Z})$. We define the $D$-intersection map
$\sigma\colon\operatorname{Div}^{0}(Y)\to V_{D}\coloneqq Z_{0}(\operatorname{\mathcal{D}})/\pi^{*}Z_{0}(\operatorname{Spec}(\mathbb{Z}))\otimes_{\mathbb{Z}}\mathbb{Q},\quad F\mapsto\Psi(F).\operatorname{\mathcal{D}}.$
Here we have $\Psi(F)=\mathcal{F}+\Phi(F)$ with $\mathcal{F}$ the extension of $F$ to a horizontal divisor on $\mathcal{X}$ obtained by taking the closure of points, $\Phi(F)$ is a sum of vertical $\mathbb{Q}$-divisors on $\mathcal{X}$ such that $\Psi(F)$ has $\ell$-intersection number 0 with all vertical divisors above $\ell$ for all primes $\ell$, and
$\Psi(F).\operatorname{\mathcal{D}}=\sum_{x}i_{x}(\Psi(F),\operatorname{\mathcal{D}})[x]$
denotes the intersection cycle of the divisors $\Psi(F)$ and $\operatorname{\mathcal{D}}$ on $\mathcal{X}$. We show that $\sigma$ descends to $J_{Y}(\mathbb{Q})$ and is independent of the choice of regular model $\mathcal{X}$. The map $\sigma$ naturally decomposes as a direct sum of local $D$-intersection maps $\sigma_{\ell}\colon J_{Y}(\mathbb{Q}_{\ell})\to V_{D,\ell}$. The construction is inspired by the Arakelov-theoretic interpretation of local Néron symbols due to Faltings and Hrijlac [Fal84, Hri85] which appear as the local components of the Néron–Tate height and of the $p$-adic Coleman–Gross height at places away from $p$ [CG89].
In the local setting, we show that the $D$-intersection map provides a $\mathbb{Q}$-form of the $p$-adic étale Kummer map of the generalised Jacobian (˜2.32). This can be seen as an “independence of pp” result and suggests an interpretation of $\sigma_{\ell}$ as a motivic Kummer map for $J_{Y}$, which may be of independent interest.
In a sense, the intersection cycle $\sigma(P-P_{0})$ records where and how badly the points $P$ and $P_{0}$ reduce to the cusps. If we restrict to $S$-integral points $P$ of a fixed reduction type, we can show that the composition $\sigma_{\ell}\circ\operatorname{AJ}_{P_{0}}\colon\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}\to V_{D,\ell}$ lands in a certain subset $\mathfrak{S}_{\ell}(P_{0},\Sigma_{\ell})\subseteq V_{D,\ell}$, which is either a point or a translate of a rank $1$ subgroup (˜2.27). Putting things together globally, we show in ˜3.13 that the Abel–Jacobi image of $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ lies in the global Selmer set
$\operatorname{Sel}(P_{0},\Sigma)\coloneqq\sigma^{-1}\left(\prod_{\ell}\mathfrak{S}_{\ell}(P_{0},\Sigma_{\ell})\right)\subseteq J_{Y}(\mathbb{Q}).$
By determining the rank of the kernel of $\sigma$ (˜3.7), we show that $\operatorname{Sel}(P_{0},\Sigma)$ is a translate of a subgroup of a certain rank (˜3.14). Hypothesis (1.2) then precisely ensures that this rank is strictly less than $\dim_{\mathbb{Q}_{p}}\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))=g+n-1$.
1.3. Structure of the paper
In Section˜2, we define the $D$-intersection map $\sigma_{\mathfrak{m}}$ in a local setting, namely for affine curves over discretely valued fields. We introduce $D$-transversal models, integral and rational reduction types $\Sigma_{\mathfrak{m}}$ and show that the composition $\sigma_{\mathfrak{m}}\circ\operatorname{AJ}_{P_{0}}$ restricted to points of a given reduction type lands in the subset $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$. We also compare the $D$-intersection map with the local Kummer map. In Section˜3, we define the global $D$-intersection map $\sigma$ for an affine curve over a number field, and we determine the rank of $\ker(\sigma)$. We define $S$-integral reduction types $\Sigma$ and the global Selmer set $\operatorname{Sel}(P_{0},\Sigma)$. We then show that the Abel–Jacobi images of $S$-integral points of reduction type $\Sigma$ are contained in $\operatorname{Sel}(P_{0},\Sigma)$ and bound the rank of $\operatorname{Sel}(P_{0},\Sigma)$. In Section˜4, we summarise the theory of $p$-adic abelian integrals of log differentials. In Section˜5 we prove ˜A and describe an $S$-integral version of Siksek’s Restriction of Scalars Chabauty [Sik13]. We then proceed to prove ˜C in Section˜6, and end this paper in Section˜7 with three applications of our methods to even degree hyperelliptic curves: recovering the Linear Quadratic Chabauty method of [GM25], an improved bound on the number of integral points of such curves, and the complete determination of the $\mathbb{Z}[\zeta_{3}]$-points of a curve of genus 2 and Mordell–Weil rank 2.
Acknowledgements
We thank Jennifer Balakrishnan, Alex Betts, Stevan Gajović, Sachi Hashimoto, Enis Kaya, Bjorn Poonen, Pim Spelier, and Elie Studnia for discussions on this paper. We thank Steffen Müller for helpful dicussions and comments on a previous version of this paper. The first author is supported by a Walter Benjamin Scholarship from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project number LE 5634/1, and also acknowledges support from the DFG through TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124. The second author is supported by a Minerva Fellowship of the Minerva Stiftung Gesellschaft für die Forschung mbH and also acknowledges support through a guest postdoc fellowship at the Max Planck Institute for Mathematics in Bonn and through an NWO Grant, project number VI.Vidi.192.106.
2. Local theory: The $D$-intersection map
We recall the definition of the generalised Jacobian and analyse the image of the Abel–Jacobi map for an affine curve over a discretely valued field. The key innovation is the construction of the so-called $D$-intersection map from $J_{Y}(K)$ to a finite-dimensional $\mathbb{Q}$-vector space $V_{D,\mathfrak{m}}$ in Section˜2.3. This map is used to constrain the image of integral and rational points of fixed reduction type under the Abel–Jacobi map.
2.1. The generalised Jacobian
Let $K$ be a field, let $X/K$ be a smooth, projective curve, let $D\neq\emptyset$ be a finite set of closed points of $X$, and set $Y=X\smallsetminus D$. The Jacobian of $X$ is denoted by $J$. The generalised Jacobian $J_{Y}$ of $Y$ [Ser88, Ch. V] is a semi-abelian variety over $K$ sitting in a short exact sequence
$0\to T_{D}\to J_{Y}\to J\to 0$ (2.1)
with the Jacobian of $X$ on the right and a torus $T_{D}$ on the left. The toric part is given by the Weil restriction $T_{D}=(\operatorname{Res}_{D/K}\mathbb{G}_{m,D})/\mathbb{G}_{m}$, whose $K$-points are given by
$T_{D}(K)=\mathcal{O}(D)^{\times}/K^{\times}=\bigl(\prod_{Q\in|D|}k(Q)^{\times}\bigr)\bigm/K^{\times}.$
Here $k(Q)$ denotes the residue field of the cusp $Q$ and the group $K^{\times}$ is embedded diagonally. The $K$-points of the generalised Jacobian $J_{Y}$ have the following interpretation: two divisors $E,E^{\prime}$ on $Y$ are called $D$-equivalent if there is a rational function $f\in k(X)^{\times}$ with $f|_{D}=1$ such that $E-E^{\prime}=\mathrm{div}(f)$. Now $J_{Y}(K)$ is the group of $D$-equivalence classes of divisors of degree zero on $Y$:
$J_{Y}(K)=\frac{\operatorname{Div}^{0}(Y)}{\left\{\mathrm{div}(f)\bigm|f\in k(X)^{\times},f|_{D}=1\right\}}.$
The map $J_{Y}\to J$ in the exact sequence (2.1) is given by viewing degree zero divisors on $Y$ as divisors on $X$ and replacing $D$-equivalence with the finer notion of linear equivalence. The inclusion $T_{D}\to J_{Y}$ is given on $K$-points as follows: given $a=(a_{Q})_{Q}\in\prod_{Q\in|D|}k(Q)^{\times}$, let $f$ be any rational function on $X$ realising the values $f(Q)=a_{Q}$ for $Q\in|D|$. Such a function exists by approximation. Then $a\mapsto\mathrm{div}(f)$ in $J_{Y}(K)$, which is well-defined and descends to $T_{D}(K)$. Given a $K$-rational point $P_{0}\in Y(K)$, we have the Abel–Jacobi map
$\operatorname{AJ}_{P_{0}}\colon Y\to J_{Y}.$
On $K$-points, it is given by mapping $P\in Y(K)$ to the $D$-equivalence class of $P]-[P_{0}]$.
2.2. Intersection numbers
For the remainder of this section, let $(R,\mathfrak{m},k)$ be a discrete valuation ring with field of fractions $K$, let $X$ be a smooth projective curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. In order to construct the $D$-intersection map $\sigma_{\mathfrak{m}}\colon J_{Y}(K)\to V_{D,\mathfrak{m}}$, we choose a regular model $\mathcal{X}$ of $X$ over $R$, i.e. an integral, regular, projective and flat $R$-scheme with an isomorphism $\mathcal{X}_{K}\cong X$ [Liu06, Definition 10.1.1]. We will show later that the map $\sigma_{\mathfrak{m}}$ does not depend on the chosen model. Recall from [Lan88, §III.2] the definition of intersection numbers on the arithmetic surface $\mathcal{X}$. If $E,F$ are effective divisors on $\mathcal{X}$ without common component and $x$ is a closed point of $\mathcal{X}_{\mathfrak{m}}$, the intersection number of $E$ and $F$ at $x$ is defined as
$i_{x}(E,F)\coloneqq\mathrm{length}_{\mathcal{O}_{\mathcal{X},x}}(\mathcal{O}_{\mathcal{X},x}/(f,g)),$
where $f$ and $g$ are defining equations for $E$ and $F$ in a neighbourhood of $x$. The symbol $i_{x}$ is bilinear, hence can be extended to not necessarily effective divisors and to $\mathbb{Q}$-divisors as long as they have no common component in their support. The intersection cycle of such divisors is defined as the zero-cycle
$E.F\coloneqq\sum_{x}i_{x}(E,F)[x],$
on $\mathcal{X}_{\mathfrak{m}}$. We also have the $\mathfrak{m}$-intersection number
$i_{\mathfrak{m}}(E,F)\coloneqq E\cdot F\coloneqq\deg_{k}(E.F)=\sum_{x}i_{x}(E,F)[k(x):k],$ (2.2)
where the sum is over all closed points of $\mathcal{X}_{\mathfrak{m}}$. By replacing a divisor with a linearly equivalent one, one can also define the self-intersection number $i_{\mathfrak{m}}(E,E)$ of a vertical prime divisor $E$. Hence, the symbol $i_{\mathfrak{m}}$ extends to all pairs of divisors on $\mathcal{X}$ with no horizontal common component [Lan88, §III.3].
Definition 2.1.
Let $\tilde{\operatorname{\mathcal{D}}}\to\operatorname{Spec}(R)$ denote the normalisation of $\operatorname{Spec}(R)$ in $\mathcal{O}(D)$. Concretely, if $D=\{Q_{1},\ldots,Q_{r}\}$ and if $R_{i}$ is the integral closure of $R$ in the residue field $k(Q_{i})$, then
$\tilde{\operatorname{\mathcal{D}}}=\coprod_{i=1}^{r}\operatorname{Spec}(R_{i}).$
For $Q\in|D|$, we write $\tilde{\mathcal{Q}}$ for the normalisation of $\operatorname{Spec}(R)$ in $k(Q)$.
We need a generalisation of intersection numbers that allows us to “intersect” divisors on $\mathcal{X}$ with $\tilde{\operatorname{\mathcal{D}}}$. By the valuative criterion of properness, the inclusion $D\hookrightarrow\mathcal{X}$ extends to a morphism $\nu\colon\tilde{\operatorname{\mathcal{D}}}\to\mathcal{X}$. Its image is the closure $\operatorname{\mathcal{D}}$ of $D$ in $\mathcal{X}$. If $\operatorname{\mathcal{D}}$ is normal, i.e. the Zariski closures of the points $Q\in|D|$ in $\mathcal{X}$ are normal and pairwise disjoint, then $\nu$ is just the inclusion of the horizontal divisor $\operatorname{\mathcal{D}}\hookrightarrow\mathcal{X}$. In general, however, $\nu$ is not a closed immersion. Nevertheless, by a slight abuse of notation, for any closed point $x\in\tilde{\operatorname{\mathcal{D}}}$ and effective divisor $E$ on $\mathcal{X}$ without common component with $\operatorname{\mathcal{D}}$, we write
$i_{x}(E,\tilde{\operatorname{\mathcal{D}}})\coloneqq\operatorname{ord}_{x}(\nu^{*}f)$ (2.3)
with $f$ a defining equation for $E$ in a neighbourhood of $\nu(x)$. From [Lan88, Lemma III.2.2] it follows that this agrees with the ordinary intersection number on $\mathcal{X}$ when $\tilde{\operatorname{\mathcal{D}}}=\operatorname{\mathcal{D}}$. Again, the definition extends to not necessarily effective divisors and to $\mathbb{Q}$-divisors $E$. We also have the intersection cycle
[…]
with $x$ running over all closed points of $\tilde{\operatorname{\mathcal{D}}}$.
2.3. The $D$-intersection map
[…]
$\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{m}}]\coloneqq\sum_{x}e_{x}[x]$
[…]
Definition 2.2.
Define the $\mathbb{Q}$-vector space $V_{D,\mathfrak{m}}$ by
$V_{D,\mathfrak{m}}\coloneqq Z_{0}(\tilde{\operatorname{\mathcal{D}}})/[\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{m}}]\otimes_{\mathbb{Z}}\mathbb{Q}.$
We first construct $\sigma_{\mathfrak{m}}$ on the level of degree-0 divisors. Given $F\in\operatorname{Div}^{0}(Y)$, by [Lan88, Theorem III.3.6], there exists a vertical $\mathbb{Q}$-divisor $\Phi(F)$ on $\mathcal{X}$ such that $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$ has $\mathfrak{m}$-intersection number zero with every vertical divisor. Here, $\mathcal{F}$ denotes the extension of $F$ to a horizontal divisor on $\mathcal{X}$ obtained by taking Zariski closures of points. Now we define
$\sigma_{\mathfrak{m}}(F)\coloneqq\Psi(F).\tilde{\operatorname{\mathcal{D}}}=\sum_{x}i_{x}(\Psi(F),\tilde{\operatorname{\mathcal{D}}})[x]\in V_{D,\mathfrak{m}},$ (2.4)
where $x$ runs over all closed points of $\tilde{\operatorname{\mathcal{D}}}$. The $\mathbb{Q}$-divisor $\Phi(F)$ is uniquely determined up to adding $\mathbb{Q}$-multiples of the complete special fibre $\mathcal{X}_{\mathfrak{m}}$, therefore $\sigma_{\mathfrak{m}}(F)=\Psi(F).\tilde{\operatorname{\mathcal{D}}}$ is well-defined modulo $\mathcal{X}_{\mathfrak{m}}.\tilde{\operatorname{\mathcal{D}}}=[\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{m}}]$, i.e. as an element of $V_{D,\mathfrak{m}}$. The bilinearity of the intersection pairing implies that $\sigma_{\mathfrak{m}}$ is a homomorphism. On principal divisors, $\sigma_{\mathfrak{m}}$ can be computed in terms of evaluation at the cusps and taking orders:
Lemma 2.3.
Let $F=\mathrm{div}(f)$ be a principal divisor with support in $Y$. Then we have
$\sigma_{\mathfrak{m}}(F)=\sum_{Q\in|D|}\sum_{x\in|\tilde{\mathcal{Q}}_{\mathfrak{m}}|}\operatorname{ord}_{x}(f(Q))[x].$
Here, $f(Q)\in k(Q)^{\times}$ denotes the evaluation of $f$ at the cusp $Q$ and $x$ runs over the points of the special fibre of the normalisation $\tilde{\mathcal{Q}}$ of $R$ in $k(Q)$.
Proof.
Let $\mathrm{div}_{\mathcal{X}}(f)$ be the divisor on $\mathcal{X}$ defined by the rational function $f$. It has horizontal part $\mathcal{F}$ and by [Lan88, Theorem III.3.1] it has $\mathfrak{m}$-intersection number zero with all vertical divisors, so we can take $\Psi(F)\coloneqq\mathrm{div}_{\mathcal{X}}(f)$ in the definition of $\sigma_{\mathfrak{m}}(F)$. By definition of the intersection numbers and (2.3) this implies
$\sigma(F)=\mathrm{div}_{\mathcal{X}}(f).\tilde{\operatorname{\mathcal{D}}}=\sum_{x\in|\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{m}}|}\operatorname{ord}_{x}(\nu^{*}f)[x],$
and the claim follows by splitting the sum over the components of $\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{m}}=\coprod_{Q\in|D|}\tilde{\mathcal{Q}}_{\mathfrak{m}}$. ∎
Lemma 2.4.
The map $\sigma_{\mathfrak{m}}\colon J_{Y}(K)\to V_{D,\mathfrak{m}}$ defined by (2.4) is well-defined.
Proof.
To see that $\sigma_{\mathfrak{m}}$ is well-defined on $D$-equivalence classes, assume that $F=\mathrm{div}(f)$ is the principal divisor on $X$ defined by a function $f\in k(X)^{\times}$ with $f|_{D}=1$. Then we have $\operatorname{ord}_{x}(f(Q))=\operatorname{ord}_{x}(1)=0$ for all $Q\in|D|$ and $x\in|\tilde{\mathcal{Q}}_{\mathfrak{m}}|$, hence $\sigma_{\mathfrak{m}}(F)=0$ by Lemma˜2.3. ∎
Definition 2.5.
We call the homomorphism $\sigma_{\mathfrak{m}}\colon J_{Y}(K)\to V_{D,\mathfrak{m}}$, $F]\mapsto\Psi(F).\tilde{\operatorname{\mathcal{D}}}$ of Lemma˜2.4 the $D$-intersection map.
Proposition 2.6.
The $D$-intersection map $\sigma_{\mathfrak{m}}\colon J_{Y}(K)\to V_{D,\mathfrak{m}}$ is independent of the choice of regular model $\mathcal{X}$.
Proof.
Assume that $\mathcal{X}$ and $\mathcal{X}^{\prime}$ are two regular models of $X$ over $R$. We want to show that the associated homomorphisms $\sigma_{\mathfrak{m}},\sigma_{\mathfrak{m}}^{\prime}\colon J_{Y}(K)\to V_{D,\mathfrak{m}}$ agree. Since any two regular models can be dominated by a third one [Lic68, Proposition 4.2], we may assume that $\mathcal{X}$ is dominated by $\mathcal{X}^{\prime}$ via a morphism of models $\pi\colon\mathcal{X}^{\prime}\to\mathcal{X}$. Let $\nu\colon\tilde{\operatorname{\mathcal{D}}}\to\mathcal{X}$ and $\nu^{\prime}\colon\tilde{\operatorname{\mathcal{D}}}\to\mathcal{X}^{\prime}$ be the morphisms obtained by extending the inclusion of $D$ into the generic fibre of $\mathcal{X}$ and $\mathcal{X}^{\prime}$, respectively. The uniqueness statement in the valuative criterion of properness implies $\nu=\pi\circ\nu^{\prime}$. Now, let $F$ be a degree-0 divisor on $Y$, let $\mathcal{F}$ be its extension to a horizontal divisor on $\mathcal{X}$, and let $\Phi(F)$ be a vertical $\mathbb{Q}$-divisor on $\mathcal{X}$ such that $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$ has $\mathfrak{m}$-intersection number zero with every vertical divisor. We claim that we can take $\Psi^{\prime}(F)\coloneqq\pi^{*}\Psi(F)$ on $\mathcal{X}^{\prime}$ in the definition of $\sigma_{\mathfrak{m}}^{\prime}(F)$. Indeed, the horizontal part of $\pi^{*}\Psi(F)$ is the proper transform of $\mathcal{F}$, which equals the horizontal extension $\mathcal{F}^{\prime}$ of $F$ on $\mathcal{X}^{\prime}$, and for any vertical divisor $V^{\prime}$ on $\mathcal{X}^{\prime}$, we have
$i_{\mathfrak{m}}(\pi^{*}\Psi(F),V^{\prime})=i_{\mathfrak{m}}(\Psi(F),\pi_{*}V^{\prime})=0$
by the projection formula [Liu06, Theorem 9.2.12(b)] and by definition of $\Psi(F)$. Hence, for any closed point $x$ on $\tilde{\operatorname{\mathcal{D}}}$, if $f\in k(\mathcal{X})^{\times}$ defines $\Psi(F)$ on $\mathcal{X}$ near $\nu(x)$, then $\pi^{*}f\in k(\mathcal{X}^{\prime})^{\times}$ defines $\Psi^{\prime}(F)$ near $\nu^{\prime}(x)$, so we have
[…]
Consequently, $\sigma_{\mathfrak{m}}^{\prime}(F)=\Psi^{\prime}(F).\tilde{\operatorname{\mathcal{D}}}=\Psi(F).\tilde{\operatorname{\mathcal{D}}}=\sigma_{\mathfrak{m}}(F)$. ∎
2.4. Computing $\Phi(F)$
In the definition of the intersection map, given a divisor $F\in\operatorname{Div}^{0}(Y)$ with horizontal extension $\mathcal{F}$, we need to find a vertical $\mathbb{Q}$-divisor $\Phi(F)$ such that $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$ has $\mathfrak{m}$-intersection number zero with all vertical divisors. We construct $\Phi(F)$ using the intersection matrix of the special fibre.
Let $\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})$ be the group of vertical divisors on $\mathcal{X}$. We have the reduction component map
$\mathrm{cpt}_{\mathfrak{m}}\colon\operatorname{Div}(X)\to\operatorname{Div}_{\mathfrak{m}}(\mathcal{X}),\quad F\mapsto\sum_{V}i_{\mathfrak{m}}(\mathcal{F},V)[V],$ (2.5)
where the sum is over all vertical prime divisors. Let $M=(i_{\mathfrak{m}}(V,W))_{V,W}$ be the intersection matrix of the (reduced) special fibre, and let $L^{+}$ be a generalised inverse of M-M, i.e. a matrix satisfying $(-M)L^{+}(-M)=-M$. For example, $L^{+}$ can be chosen as the Moore–Penrose pseudoinverse of M-M. Denote the endomorphism of $\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})\otimes\mathbb{Q}$ represented by $L^{+}$ in the standard basis also by $L^{+}$.
Lemma 2.7.
For $F\in\operatorname{Div}^{0}(X)$, the vertical $\mathbb{Q}$-divisor $\Phi(F)\coloneqq L^{+}(\mathrm{cpt}_{\mathfrak{m}}(F))$ is such that $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$ has $\mathfrak{m}$-intersection number zero with all vertical divisors.
Proof.
The intersection matrix $M$ represents the endomorphism of $\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})\otimes\mathbb{Q}$ given by $E\mapsto\sum_{V}i_{\mathfrak{m}}(E,V)[V]$, with $V$ running over all vertical prime divisors. Let $m_{V}$ be the multiplicity of $V$ in the special fibre, so that $\mathcal{X}_{\mathfrak{m}}=\sum m_{V}V$. Define a degree function on $\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})\otimes\mathbb{Q}$ by setting $\deg(V)\coloneqq m_{V}$ and extending linearly. Then the endomorphism $M$ maps to the degree-zero subspace $\operatorname{Div}_{\mathfrak{m}}^{0}(\mathcal{X})\otimes\mathbb{Q}$:
$\deg(M(E))=\sum_{V}m_{V}i_{\mathfrak{m}}(E,V)=i_{\mathfrak{m}}(E,\mathcal{X}_{\mathfrak{m}})=0.$
Here we used that the complete special fibre $\mathcal{X}_{\mathfrak{m}}$ has $\mathfrak{m}$-intersection number zero with any vertical divisor [Lan88, Proposition III.3.2]. Since the intersection matrix $M$ has corank 1 [Lan88, Lemma III.3.4], the image of the endomorphism $M$ is precisely the degree-zero subspace $\operatorname{Div}_{\mathfrak{m}}^{0}(\mathcal{X})\otimes\mathbb{Q}$. (In loc. cit., the intersection matrix is defined as $(m_{V}m_{W}i_{\mathfrak{m}}(V,W))_{V,W}$ but rescaling the rows and columns does not change the corank.) The reduction component map $\mathrm{cpt}_{\mathfrak{m}}$ defined in (2.5) preserves the degree:
$\deg(\mathrm{cpt}_{\mathfrak{m}}(F))=\sum_{V}m_{V}i_{\mathfrak{m}}(\mathcal{F},V)=i_{\mathfrak{m}}(\mathcal{F},\mathcal{X}_{\mathfrak{m}})=\deg(F).$
The last equality follows from the fact that the horizontal extension of a point $Q$ of the generic fibre intersects the special fibre in $\deg(Q)=[k(Q):K]$ many points when counted with multiplicities [Lan88, Proposition III.2.5].
Now assume that $F$ is a divisor on $X$ of degree 0. Then by the above, its reduction $\mathrm{cpt}_{\mathfrak{m}}(F)$ is contained in the degree-zero subspace of $\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})$, hence in the image of the endomorphism $M$. Since $L^{+}$ is a generalised inverse of M-M, the $\mathbb{Q}$-divisor $\Phi(F)\coloneqq L^{+}(\mathrm{cpt}_{\mathfrak{m}}(F))$ satisfies $M(\Phi(F))=-\mathrm{cpt}_{\mathfrak{m}}(F)$. Unfolding the definitions of $M$ and $\mathrm{cpt}_{\mathfrak{m}}$, this means that $i_{\mathfrak{m}}(\Phi(F),V)=-i_{\mathfrak{m}}(\mathcal{F},V)$ for all vertical prime divisors $V$. Hence, $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$ has $\mathfrak{m}$-intersection number zero with all vertical divisors as claimed. ∎
Lemma 2.8.
For $F\in\operatorname{Div}^{0}(X)$, the vertical $\mathbb{Q}$-divisor $\Phi(F)\in\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})/[\mathcal{X}_{\mathfrak{m}}]\otimes\mathbb{Q}$ for which $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$ has $\mathfrak{m}$-intersection number zero with all vertical divisors depends only on the image of $F$ under the reduction component map $\mathrm{cpt}_{\mathfrak{m}}$ defined in (2.5).
Proof.
This is an immediate consequence of Lemma˜2.7. ∎
In light of Lemma˜2.8, we also write $\Phi(\mathrm{cpt}_{\mathfrak{m}}(F))$ instead of $\Phi(F)$. More generally:
Definition 2.9.
Given a vertical $\mathbb{Q}$-divisor $E\in\operatorname{Div}_{\mathfrak{m}}^{0}(\mathcal{X})$ of degree zero (with respect to the degree function in the proof of Lemma˜2.7), define $\Phi(E)\coloneqq L^{+}(E)\in\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})\otimes\mathbb{Q}$, so that $\Phi(E)$ is a vertical $\mathbb{Q}$-divisor satisfying $i_{\mathfrak{m}}(\Phi(E),V)=-\mathrm{mult}_{V}(E)$ for all vertical divisors $V$. Note that $\Phi(E)$ is well-defined up to adding $\mathbb{Q}$-multiples of $\mathcal{X}_{\mathfrak{m}}$.
Another consequence of Lemma˜2.7 is that the denominators of the coefficients of the $\mathbb{Q}$-divisors $\Phi(F)$ are bounded. Namely, let $a$ be the least common denominator of all entries of $L^{+}$. For an abelian group $B$ let $B_{\operatorname{tors}}$ be its torsion subgroup and $B_{\operatorname{tf}}\coloneqq B/B_{\operatorname{tors}}$. With $\Phi$ defined as in Lemma˜2.7, the $D$-intersection map $\sigma_{\mathfrak{m}}$ in fact takes values in the finitely generated free abelian group
$A_{D,\mathfrak{m}}\coloneqq\frac{1}{a}\left(Z_{0}(\tilde{\operatorname{\mathcal{D}}})/[\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{m}}]\right)_{\operatorname{tf}}\subseteq V_{D,\mathfrak{m}}.$ (2.6)
The $\mathbb{Z}$-structure $A_{D,\mathfrak{m}}$ of the $\mathbb{Q}$-vector space $V_{D,\mathfrak{m}}$ will be useful in §2.8.
2.5. Integral reduction types
Let $\mathcal{X}$ be a regular model of $X$ over $R$. Let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. Any $K$-rational point $P\in Y(K)$ extends uniquely to an $R$-valued point of $\mathcal{X}$ by properness of $\mathcal{X}$. Reducing this point modulo $\mathfrak{m}$ yields a $k$-point of the special fibre. This point is contained in the smooth locus $\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}$ of $\mathcal{X}_{\mathfrak{m}}$, see [LT16, Lemma 3.1].
Definition 2.10.
The composition
${\mathrm{red}}_{\mathfrak{m}}\colon Y(K)\hookrightarrow X(K)=\mathcal{X}(R)\to\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}(k)$
sending a $K$-rational point of $Y$ to its reduction in the special fibre of $\mathcal{X}$ is called the mod-$\mathfrak{m}$ reduction map.
A $K$-point of $Y$ belongs to the subset of $R$-integral points $\mathcal{Y}(R)\subseteq Y(K)$ if and only if its mod-$\mathfrak{m}$ reduction does not lie on $\operatorname{\mathcal{D}}_{\mathfrak{m}}$. The reduction type of such a point remembers the component of the special fibre onto which it reduces.
Definition 2.11.
An integral mod-$\mathfrak{m}$ reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$ is an irreducible (equivalently: connected) component of $\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}_{\mathfrak{m}}$ that contains at least one smooth $k$-point. The set of all integral mod-$\mathfrak{m}$ reduction types is denoted by $\mathcal{C}_{\mathfrak{m}}^{0}$. We say that an $R$-integral point $P$ of $\mathcal{Y}$ has reduction type $\Sigma_{\mathfrak{m}}$ if its mod-$\mathfrak{m}$ reduction lies on $\Sigma_{\mathfrak{m}}$. The set of all $R$-integral points of reduction type $\Sigma_{\mathfrak{m}}$ is denoted by $\mathcal{Y}(R)_{\Sigma_{\mathfrak{m}}}$.
Integral reduction types define a partition of $\mathcal{Y}(R)$:
$\mathcal{Y}(R)=\coprod_{\Sigma_{\mathfrak{m}}\in\mathcal{C}^{0}_{\mathfrak{m}}}\mathcal{Y}(R)_{\Sigma_{\mathfrak{m}}}.$
Remark 2.12.
Our notion of integral reduction type differs slightly from the reduction types considered in [Bet23, §6.1.1]. By only admitting components of the smooth locus $\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}$ of the special fibre we exclude all components of multiplicity larger than one. The reason for this is that only components of the smooth locus can occur as reduction types of $R$-integral points.
The reduction type of a point puts strong constraints on its image under the Abel–Jacobi map.
Proposition 2.13.
Let $P_{0}\in Y(K)$. Then the image of an $R$-integral point of $\mathcal{Y}$ under the composition $\sigma_{\mathfrak{m}}\circ\operatorname{AJ}_{P_{0}}\colon\mathcal{Y}(R)\to J_{Y}(K)\to V_{D,\mathfrak{m}}$ depends only on its reduction type. More precisely, if $P\in\mathcal{Y}(R)$ has reduction type $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{0}$ then
$\sigma_{\mathfrak{m}}(P-P_{0})=-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(\Sigma_{\mathfrak{m}}-\mathrm{cpt}_{\mathfrak{m}}(P_{0})).\tilde{\operatorname{\mathcal{D}}}$. (2.7)
Proof.
Let $P$ be an $R$-integral point of reduction type $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{0}$. Let $\mathcal{P}$ and $\mathcal{P}_{0}$ be the horizontal divisors on $\mathcal{X}$ given by taking the Zariski closure of $P$ and $P_{0}$, respectively. We calculate
$\sigma_{\mathfrak{m}}(P-P_{0})=\Psi(P-P_{0}).\tilde{\operatorname{\mathcal{D}}}=\mathcal{P}.\tilde{\operatorname{\mathcal{D}}}-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(P-P_{0}).\tilde{\operatorname{\mathcal{D}}}$.
Since $P$ is an $R$-integral point, its reduction does not lie on $\operatorname{\mathcal{D}}$, so the first summand vanishes. The second summand is independent of $P$. The third summand, by Lemma˜2.8, depends on $P$ only through its reduction component $\mathrm{cpt}_{\mathfrak{m}}(P)$. By assumption, $P$ reduces onto the component given by the reduction type $\Sigma_{\mathfrak{m}}$, and the $\mathfrak{m}$-intersection number of $\mathcal{P}$ with this component equals 1 since $i_{\mathfrak{m}}(\mathcal{P},\mathcal{X}_{\mathfrak{m}})=[k(P):K]=1$, see [Lan88, Proposition 2.5]. So we have $\mathrm{cpt}_{\mathfrak{m}}(P)=\Sigma_{\mathfrak{m}}$, viewing $\Sigma_{\mathfrak{m}}$ as an element of $\operatorname{Div}_{\mathfrak{m}}(\mathcal{X})$ in the obvious way. According to Lemma˜2.7, we can take
$\Phi(P-P_{0})=L^{+}(\mathrm{cpt}_{\mathfrak{m}}(P-P_{0}))=\Phi(\Sigma_{\mathfrak{m}}-\mathrm{cpt}_{\mathfrak{m}}(P_{0})),$
which gives the claimed formula (2.7). ∎
Definition 2.14.
Let $P_{0}\in Y(K)$ be a base point and let $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{0}$ be an integral mod-$\mathfrak{m}$ reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$. We define $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$ to be the one-element subset of $V_{D,\mathfrak{m}}$ containing the right hand side of (2.7), i.e.
$\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})=\bigl\{-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(\Sigma_{\mathfrak{m}}-\mathrm{cpt}_{\mathfrak{m}}(P_{0})).\tilde{\operatorname{\mathcal{D}}}\bigr\}$. (2.8)
According to ˜2.13, we have
$(\sigma_{\mathfrak{m}}\circ\operatorname{AJ}_{P_{0}})(P)\in\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$
for all $P\in\mathcal{Y}(R)$ of reduction type $\Sigma_{\mathfrak{m}}$.
Proposition 2.15.
Let $P_{0}\in Y(K)$ be a base point and let $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{0}$ be an integral mod-$\mathfrak{m}$ reduction type. If $P_{0}$ is itself an $R$-integral point of reduction type $\Sigma_{\mathfrak{m}}$, then $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})=\{0\}$.
Proof.
If $P_{0}$ is $R$-integral then its closure $\mathcal{P}_{0}$ does not meet the closure $\operatorname{\mathcal{D}}$ of $D$ in $\mathcal{X}$, so that we have $\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}=0$. If moreover $P_{0}$ has reduction type $\Sigma_{\mathfrak{m}}$, then $\mathrm{cpt}_{\mathfrak{m}}(P_{0})=\Sigma_{\mathfrak{m}}$, hence we can choose $\Phi(\Sigma_{\mathfrak{m}}-\mathrm{cpt}_{\mathfrak{m}}(P_{0}))=\Phi(0)=0$, which implies $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})=\{0\}$ by (2.8). ∎
It can happen that the composition $\sigma_{\mathfrak{m}}\circ\operatorname{AJ}_{P_{0}}\colon\mathcal{Y}(R)\to J_{Y}(K)\to V_{D,\mathfrak{m}}$ is constant on all of $\mathcal{Y}(R)$, not only on points of a fixed reduction type. The following proposition gives a criterion for when this occurs.
Proposition 2.16.
Assume that $\operatorname{\mathcal{D}}$ intersects the special fibre $\mathcal{X}_{\mathfrak{m}}$ in only one component. Then the set $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$ is equal to $\{-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}\}$, independently of the choice of integral reduction type $\Sigma_{\mathfrak{m}}$.
Proof.
According to (2.8), $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$ is the singleton set containing an element of the form $\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+V.\tilde{\operatorname{\mathcal{D}}}$ where $V$ is a vertical $\mathbb{Q}$-divisor depending on $\Sigma_{\mathfrak{m}}$. We claim that $V.\tilde{\operatorname{\mathcal{D}}}=0$ in $V_{D,\mathfrak{m}}$ for every vertical divisor $V$. We may assume that $V$ is a prime divisor, i.e. a component of the special fibre. Let $m_{V}$ be the multiplicity of $V$ in $\mathcal{X}_{\mathfrak{m}}$. By assumption, either $\operatorname{\mathcal{D}}$ does not intersect $V$, or all points of the intersection $\operatorname{\mathcal{D}}\cap\mathcal{X}_{\mathfrak{m}}$ are contained in $V$. In the former case, we trivially have $V.\tilde{\operatorname{\mathcal{D}}}=0$, while in the latter case we have
$V.\tilde{\operatorname{\mathcal{D}}}=\tfrac{1}{m_{V}}\mathcal{X}_{\mathfrak{m}}.\tilde{\operatorname{\mathcal{D}}}=\tfrac{1}{m_{V}}[\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{m}}],$
but in $V_{D,\mathfrak{m}}$ this is zero as well. ∎
2.6. $D$-transversal models
Let $\mathcal{X}$ be a regular model of $X$ over $R$, and let $\operatorname{\mathcal{D}}$ and $\mathcal{Y}$ be as before. We also want to partition the set $Y(K)$, not just $\mathcal{Y}(R)$, by reduction type. But $P\in Y(K)$ may reduce to the reduction of a cusp. If this happens, we want to record the cusp to which $P$ reduces, so we need the model $\mathcal{X}$ to be $D$-transversal:
Definition 2.17.
  1. (1)
    Let $x$ be a point in $(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$. We say that $\mathcal{X}$ is $D$-transversal at $x$ if $i_{x}(\mathcal{X}_{\mathfrak{m}},\operatorname{\mathcal{D}})=1$.
  2. (2)
    We say that $\mathcal{X}$ or $(\mathcal{X},\operatorname{\mathcal{D}})$ is $D$-transversal if it is $D$-transversal at all $x\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$.
By [Liu06, Lemma 9.1.8(b)], if the model $\mathcal{X}$ is $D$-transversal at $x\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$ then $\operatorname{\mathcal{D}}$ is regular (equivalently, normal) at $x$. In particular, no two components of $\operatorname{\mathcal{D}}$ intersect at $x$, so $x$ lies in the closure of a unique point $Q\in|D|$. Note that $\operatorname{\mathcal{D}}$ may be non-normal at points where it intersects the special fibre with higher multiplicity (e.g., singular points of $\mathcal{X}_{\mathfrak{m}}$) or at points which are not $k$-rational. However, $K$-points of $Y$ reduce to smooth $k$-rational points of the special fibre, so a transversality condition at these points is all we need.
Remark 2.18.
[Bet18, Appendix B] defines a regular normal crossings (RNC) model $\mathcal{X}$ as a regular model for which $\operatorname{\mathcal{D}}+\mathcal{X}_{\mathfrak{m}}$ is a normal crossings divisor on $\mathcal{X}$. Every RNC model is $D$-transversal, but the converse is not true.
Proposition 2.19.
If $R$ is excellent, then any regular model $(\mathcal{X},\operatorname{\mathcal{D}})$ is dominated by a $D$-transversal regular model.
Proof.
As $R$ is excellent, by [Liu06, Theorem 9.2.26] successive blow-ups of $\mathcal{X}$ in closed points [^1] As we only need $\operatorname{\mathcal{D}}$ to be regular at all $x\in\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}(k)$, blow-ups in those points $x\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$ with $i_{x}(\mathcal{X}_{\mathfrak{m}},\operatorname{\mathcal{D}})\geq 2$ should suffice, but we did not check whether the analogue of [Liu06, Lemma 9.2.32] is true. yield a regular model $\mathcal{X}^{\prime}$ of $X$ in which the closure $\operatorname{\mathcal{D}}^{\prime}$ of $D$ is normal. Now we blow up again in every $x\in({\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$ such that $i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})\geq 2$ to obtain the model $\mathcal{X}^{\prime\prime}$. We claim that $\mathcal{X}^{\prime\prime}$ is a $D$-transversal regular model of $X$. Indeed, away from those $x$ the model $\mathcal{X}^{\prime}$ is already $D$-transversal. For $x\in({\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$, let $Q$ be the unique cusp such that $x$ lies in the closure $\mathcal{Q}^{\prime}$ of $Q$. If $2\leq i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})=i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\mathcal{Q}^{\prime})$, let $E$ denote the exceptional divisor on $\mathcal{X}^{\prime\prime}$ of the blow-up in $x$. Then $\mathcal{Q}^{\prime\prime}$ intersects $\mathcal{X}^{\prime\prime}_{\mathfrak{m}}$ in the intersection point of $E$ with the strict transform of ${\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}$, which is not a point in ${\mathcal{X}^{\prime\prime}_{\mathfrak{m}}}^{\mathrm{sm}}$, so there is no $D$-transversality left to check. ∎
Definition 2.20.
A regular model $(\mathcal{X},\operatorname{\mathcal{D}})$ of $(X,D)$ over $R$ is the minimal regular $D$-transversal model if it is dominated by any other such model.
It is clear that a minimal regular $D$-transversal model, if it exists, is unique up to unique isomorphism.
Proposition 2.21.
If $Y$ is hyperbolic and $R$ is excellent, then a minimal regular $D$-transversal model exists.
For the proof, we need a criterion characterising exceptional divisors on a $D$-transversal model whose contraction is again $D$-transversal.
Definition 2.22.
Let $\mathcal{X}$ be a $D$-transversal model and let $E$ be an exceptional divisor on $\mathcal{X}$. We say $E$ is $D$-exceptional if the contraction of $E$ is again a $D$-transversal model.
Lemma 2.23.
Let $\mathcal{X}$ be a $D$-transversal model.
  1. (1)
    The blow-up of $\mathcal{X}$ in a closed point $x$ is again $D$-transversal.
  2. (2)
    Let $E$ be an exceptional divisor on $\mathcal{X}$. Then $E$ is $D$-exceptional if and only if (at least) one of the following conditions is satisfied:
    1. (a)
      $\mathrm{H}^{0}(E,\mathcal{O}_{E})\neq k$.
    2. (b)
      $E$ has multiplicity $\geq 2$ as a component of $\mathcal{X}_{\mathfrak{m}}$.
    3. (c)
      $E$ and $\operatorname{\mathcal{D}}$ do not intersect.
    4. (d)
      $E$ meets $\operatorname{\mathcal{D}}$ in a unique closed point $x$, which is $k$-rational and lies in $\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}$.
Proof.
(1) Let $\pi\colon\mathcal{X}^{\prime}\to\mathcal{X}$ be the blow-up of $\mathcal{X}$ in $x$, let $E=\pi^{-1}(x)$ denote the exceptional locus of $\pi$ and let $\operatorname{\mathcal{D}}^{\prime}$ be the closure of $D$ in $\mathcal{X}^{\prime}$. As $\pi$ is an isomorphism above $\mathcal{X}\smallsetminus\{x\}$, we only need to check that $\mathcal{X}^{\prime}$ is $D$-transversal at points in $E$. Now if $x\not\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$ then $E$ is not defined over $k$ or does not have multiplicity $1$ as a component of $\mathcal{X}^{\prime}_{\mathfrak{m}}$ or does not intersect $\operatorname{\mathcal{D}}^{\prime}$, and in all cases $\mathcal{X}^{\prime}$ is automatically $D$-transversal. So assume $x\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$. The projection formula [Lan88, Theorem III.4.1] applied to the divisors $\mathcal{X}_{\mathfrak{m}}$ on $\mathcal{X}$ and $\operatorname{\mathcal{D}}^{\prime}$ on $\mathcal{X}^{\prime}$ yields
$\pi_{*}(\pi^{*}\mathcal{X}_{\mathfrak{m}}.\operatorname{\mathcal{D}}^{\prime})=\mathcal{X}_{\mathfrak{m}}.\pi_{*}(\operatorname{\mathcal{D}}^{\prime}).$
Using that $\pi^{*}(\mathcal{X}_{\mathfrak{m}})=\mathcal{X}_{\mathfrak{m}}^{\prime}$ and $\pi_{*}(\operatorname{\mathcal{D}}^{\prime})=\operatorname{\mathcal{D}}$, comparing coefficients of $[x]$ yields
$\sum_{x^{\prime}\in E}i_{x^{\prime}}(\mathcal{X}_{\mathfrak{m}}^{\prime},\operatorname{\mathcal{D}}^{\prime})[k(x^{\prime}):k(x)]=i_{x}(\mathcal{X}_{\mathfrak{m}},\operatorname{\mathcal{D}}).$ (2.9)
As $\mathcal{X}$ is $D$-transversal at $x$, the right hand side is equal to $1$, which implies that there is exactly one closed point $x^{\prime}$ where $E$ and $\operatorname{\mathcal{D}}^{\prime}$ intersect, and this point is $k$-rational and satisfies $i_{x^{\prime}}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})=1$, i.e. $\mathcal{X}^{\prime}$ is $D$-transversal at $x^{\prime}$.
(2) Let $\pi\colon\mathcal{X}\to\mathcal{X}^{\prime}$ be the contraction of $E$, let $x^{\prime}\coloneqq\pi(E)$, and let $\operatorname{\mathcal{D}}^{\prime}$ be the closure of $D$ in $\mathcal{X}^{\prime}$. As $\pi$ is an isomorphism above $\mathcal{X}^{\prime}\smallsetminus\{x^{\prime}\}$, we see that $\mathcal{X}^{\prime}$ is automatically $D$-transversal except possibly at $x^{\prime}$. Equation (2.9) applied to $\pi\colon\mathcal{X}\to\mathcal{X}^{\prime}$ yields
$\sum_{x\in E}i_{x}(\mathcal{X}_{\mathfrak{m}},\operatorname{\mathcal{D}})[k(x):k(x^{\prime})]=i_{x^{\prime}}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime}).$ (2.10)
Now in cases 2(a), 2(b), or 2(c), we have $x^{\prime}\not\in({\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$ and hence $\mathcal{X}^{\prime}$ is automatically $D$-transversal. In case 2(d), we have $x^{\prime}\in({\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$, so $\mathcal{X}$ is $D$-transversal at $x$ by assumption and $i_{x^{\prime}}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})=1$ by (2.10), as desired.
Conversely, suppose that $\mathcal{X}^{\prime}$ is $D$-transversal and assume we are not in cases 2(a), 2(b), or 2(c). That is, we assume that $\mathrm{H}^{0}(E,\mathcal{O}_{E})=k(x^{\prime})=k$, that $E$ has multiplicity $1$ in $\mathcal{X}_{\mathfrak{m}}$, which is equivalent by [Liu06, Proposition 9.2.23] to $\mathcal{X}^{\prime}_{\mathfrak{m}}$ being regular at $x^{\prime}$, and that $x^{\prime}$ lies on $\operatorname{\mathcal{D}}^{\prime}$. Then $x^{\prime}\in({\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$, so $i_{x^{\prime}}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})=1$ by $D$-transversality, and (2.10) implies condition 2(d). ∎
Proof of ˜2.21.
˜2.19 ensures the existence of a $D$-transversal regular model. By repeatedly contracting all $D$-exceptional divisors, we obtain a relatively minimal $D$-transversal model, i.e. one that does not dominate any other such model. We show that any two relatively minimal $D$-transversal models $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are isomorphic, and therefore minimal such models. Assume the contrary. As $\mathcal{X}_{1}$ is relatively minimal, the birational map $f\colon\mathcal{X}_{1}\dashrightarrow\mathcal{X}_{2}$ is not defined everywhere, say not at the closed point $x_{0}$. We now refine the argument of the corrected version of [Liu06, Lemma 9.3.20], published as an erratum [LiuErr] on the author’s webpage. Namely, there exists a morphism $g\colon\widetilde{\mathcal{Z}}\to\mathcal{X}_{2}$ that is a finite sequence of blow-ups in closed points
$g\colon\widetilde{\mathcal{Z}}=\mathcal{Z}_{n}\to\ldots\to\mathcal{Z}_{1}\to\mathcal{Z}_{0}=\mathcal{X}_{2}$
and a morphism $h\colon\widetilde{\mathcal{Z}}\to\mathcal{X}_{1}$ such that the diagram
$\widetilde{\mathcal{Z}}$$\mathcal{X}_{2}$$\mathcal{X}_{1}$$\scriptstyle{g}$$\scriptstyle{h}$$\scriptstyle{f^{-1}}$
commutes. Let $E$ be an exceptional divisor on $\widetilde{\mathcal{Z}}$ contained in $h^{-1}(x_{0})$ and let $\Gamma_{i}\subseteq\mathcal{Z}_{i}$ denote the exceptional divisor of $\mathcal{Z}_{i}\to\mathcal{Z}_{i-1}$. By Lemma˜2.23, the model $\widetilde{\mathcal{Z}}$ is $D$-transversal and the divisors $E$ and all $\Gamma_{i}$ are $D$-exceptional. If $E$ and $\Gamma_{n}$ do not meet, the image of $E$ in $\mathcal{Z}_{n-1}$ is again $D$-exceptional and will also be denoted by $E$. Let $m\leq n$ be the smallest integer such that $\Gamma_{m}$ and $E$ are disjoint. If $m=1$, then $E$ is a $D$-exceptional divisor on $\mathcal{Z}_{0}=\mathcal{X}_{2}$, contradicting the relative minimality of $\mathcal{X}_{2}$. If $m\geq 2$, then $E$ and $\Gamma_{m-1}$ are exceptional divisors on $\mathcal{Z}_{m-1}$ such that $E\cap\Gamma_{m-1}\neq\emptyset$. Pick a point $x\in E\cap\Gamma_{m-1}$ and let $k_{1}=\mathrm{H}^{0}(\Gamma_{m-1},\mathcal{O}_{\Gamma_{m-1}})$ and $k_{2}=\mathrm{H}^{0}(E,\mathcal{O}_{E})$. Then $k_{i}\subseteq k(x)$ and
$(\Gamma_{m-1}+E)^{2}=\Gamma_{m-1}^{2}+E^{2}+2\Gamma_{m-1}\cdot E\geq-[k_{1}:k]-[k_{2}:k]+2[k(x):k]\geq 0.$
By [Liu06, Theorem 9.1.23], we have equality above, thus $k_{1}=k(x)=k_{2}$, and we have $\mathcal{Z}_{m-1,\mathfrak{m}}=i(\Gamma_{m-1}+E)$ for some integer $i\geq 1$. The adjunction formula [Liu06, Propositions 9.1.35 and 9.3.10(a)] shows $g=0$; in particular, if $X$ has genus $g\geq 1$ we have arrived at a contradiction. Assume $g=0$ from now on. As the smooth locus of $\mathcal{Z}_{m-1,\mathfrak{m}}$ cannot be empty, we have $i=1$. As $X$ is geometrically connected, so is the special fibre $\mathcal{Z}_{m-1,\mathfrak{m}}$ by [Liu06, Lemma 8.3.6(b)] and hence $k_{1}=k_{2}=k$. As $\Gamma\in\{\Gamma_{m-1},E\}$ is $D$-exceptional, Lemma˜2.23(2) shows that $\Gamma\cdot\operatorname{\mathcal{D}}\leq 1$, where $\operatorname{\mathcal{D}}$ is the closure of $D$ inside the model $\mathcal{Z}_{m-1}$. Now $Y$ is hyperbolic, hence $D(\overline{K})\geq 3$, and using [Liu06, Proposition 9.1.30] we arrive at the contradiction
$3\leq\#D(\overline{K})=\sum_{Q\in|D|}[k(Q):K]=\mathcal{Z}_{m-1,\mathfrak{m}}\cdot\operatorname{\mathcal{D}}=(\Gamma_{m-1}+E)\cdot\operatorname{\mathcal{D}}\leq 2.\qed$
Remark 2.24.
In all applications, $R$ will be the localisation of the ring of integers in a number field at a prime ideal or a completion thereof and hence excellent, so the existence of a $D$-transversal regular model is guaranteed by ˜2.21. Working with one such model is usually enough, but the minimal $D$-transversal regular model minimizes the number of possible rational reduction types, compare §2.7.
2.7. Rational reduction types
The $K$-rational points $Y(K)$ can also be partitioned by reduction type, according to their mod-$\mathfrak{m}$ reduction on a suitable regular model. Here we want to work with a $D$-transversal regular model $(\mathcal{X},\operatorname{\mathcal{D}})$, see ˜2.17. A $K$-point of $Y$, unlike an $R$-integral point, may reduce onto the special fibre of $\operatorname{\mathcal{D}}$. In this case, the reduction type of the point records to which point on $\operatorname{\mathcal{D}}_{\mathfrak{m}}$ it reduces; otherwise, it records only the component.
Definition 2.25.
  1. (a)
    Let $(\mathcal{X},\operatorname{\mathcal{D}})$ be a regular $D$-transversal model of $(X,D)$. A rational mod-$\mathfrak{m}$ reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$ is either an irreducible/connected component of $\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}_{\mathfrak{m}}$ containing at least one $k$-point, or a $k$-point on the special fibre of $\operatorname{\mathcal{D}}$ at which $\mathcal{X}_{\mathfrak{m}}$ is smooth. In the latter case, we call the reduction type cuspidal; otherwise we call it a component reduction type. The set of rational mod-$\mathfrak{m}$ reduction types is denoted by $\mathcal{C}_{\mathfrak{m}}\coloneqq\mathcal{C}_{\mathfrak{m}}^{0}\sqcup\mathcal{C}_{\mathfrak{m}}^{1}$ where
    $\mathcal{C}_{\mathfrak{m}}^{0}$ $\coloneqq\{\text{components of $\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}$ containing a $k$-point}\},$
    $\mathcal{C}_{\mathfrak{m}}^{1}$ $\coloneqq(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k).$
    For a reduction type $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}$, denote by $Y(K)_{\Sigma_{\mathfrak{m}}}$ the set of all $K$-rational points of $Y$ whose mod-$\mathfrak{m}$ reduction lies on $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{0}$, respectively, is equal to $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{1}$.
  2. (b)
    If no model of $X$ is specified and $Y$ is hyperbolic, a rational mod-$\mathfrak{m}$ reduction type for $(X,D)$ is a rational mod-$\mathfrak{m}$ reduction type for the minimal regular $D$-transversal model $(\mathcal{X}_{\min},\operatorname{\mathcal{D}}_{\min})$, which exists by ˜2.21.
For any regular $D$-transversal model $(\mathcal{X},\operatorname{\mathcal{D}})$ of $(X,D)$, the rational mod-$\mathfrak{m}$ reduction types for $(\mathcal{X},\operatorname{\mathcal{D}})$ define a partition of $Y(K)$:
$Y(K)=\coprod_{\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}}Y(K)_{\Sigma_{\mathfrak{m}}}.$
Remark 2.26.
Our rational mod-$\mathfrak{m}$ reduction types are a slight modification of the reduction types considered in [Bet23, §6.1.2]. One difference is that we only require the model to be $D$-transversal, which is weaker than requiring the divisor $\mathcal{X}_{\mathfrak{m}}+\operatorname{\mathcal{D}}$ to have normal crossings everywhere. Moreover, our $\mathcal{C}_{\mathfrak{m}}^{1}$ contains only $k$-points rather than all closed points of $\operatorname{\mathcal{D}}_{\mathfrak{m}}$, and by requiring $\mathcal{X}_{\mathfrak{m}}$ to be smooth at those points we exclude points where $\operatorname{\mathcal{D}}$ intersects the special fibre $\mathcal{X}_{\mathfrak{m}}$ in a component of higher multiplicity. Since every $K$-point of $Y$ reduces onto a $k$-point of the smooth locus of $\mathcal{X}_{\mathfrak{m}}$, our reduction types are the only ones that can occur.
For $K$-rational points, the reduction type constrains the image in $V_{D,\mathfrak{m}}$ to a translate of a subgroup of rank $\leq 1$.
Proposition 2.27.
Let $P_{0}\in Y(K)$, let $(\mathcal{X},\operatorname{\mathcal{D}})$ be a regular $D$-transversal model of $(X,D)$ and let $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}$ be a rational mod-$\mathfrak{m}$ reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$. Consider the composition
$Y(K)\overset{\operatorname{AJ}_{P_{0}}}{\longrightarrow}J_{Y}(K)\overset{\sigma_{\mathfrak{m}}}{\longrightarrow}V_{D,\mathfrak{m}}.$ (2.11)
  1. (a)
    If $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{0}$ is a component then the map (2.11) is constant on $Y(K)_{\Sigma_{\mathfrak{m}}}$. More precisely, if $P\in Y(K)$ has reduction type $\Sigma_{\mathfrak{m}}$ then
    $\sigma_{\mathfrak{m}}(P-P_{0})=-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(\Sigma_{\mathfrak{m}}-\mathrm{cpt}_{\mathfrak{m}}(P_{0})).\tilde{\operatorname{\mathcal{D}}}$. (2.12)
  2. (b)
    If $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{1}$ is a point in $(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$, then the image of $Y(K)_{\Sigma_{\mathfrak{m}}}$ under (2.11) is contained in a translate of a subgroup of $V_{D,\mathfrak{m}}$ of rank $\leq 1$. More precisely, if $P\in Y(K)$ has reduction type $\Sigma_{\mathfrak{m}}$ then
    $\sigma_{\mathfrak{m}}(P-P_{0})\in\mathbb{Z}\cdot[\Sigma_{\mathfrak{m}}]-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(\mathrm{cpt}(\Sigma_{\mathfrak{m}})-\mathrm{cpt}_{\mathfrak{m}}(P_{0})).\tilde{\operatorname{\mathcal{D}}}$. (2.13)
Proof.
In case (a), $\Sigma_{\mathfrak{m}}$ is an integral reduction type, so $Y(K)_{\Sigma_{\mathfrak{m}}}=\mathcal{Y}(R)_{\Sigma_{\mathfrak{m}}}$ and the claim follows from ˜2.13. Suppose we are in case (b). Let $P$ be a $K$-rational point of reduction type $\Sigma_{\mathfrak{m}}$, let $\mathcal{P}$ and $\mathcal{P}_{0}$ be the horizontal extension of $P$ and $P_{0}$ in $\mathcal{X}$, respectively. Then
$\sigma_{\mathfrak{m}}(P-P_{0})=\Psi(P-P_{0}).\tilde{\operatorname{\mathcal{D}}}=\mathcal{P}.\operatorname{\mathcal{D}}-\mathcal{P}_{0}.\operatorname{\mathcal{D}}+\Phi(P-P_{0}).\tilde{\operatorname{\mathcal{D}}}$.
Since $P$ reduces to $\Sigma_{\mathfrak{m}}\in\operatorname{\mathcal{D}}_{\mathfrak{m}}$ and $\mathcal{X}$ is $D$-transversal, the divisor $\mathcal{P}$ intersects $\tilde{\operatorname{\mathcal{D}}}$ precisely in the point $\Sigma_{\mathfrak{m}}$, so we have $\mathcal{P}.\tilde{\operatorname{\mathcal{D}}}\in\mathbb{Z}\cdot[\Sigma_{\mathfrak{m}}]$. For the third summand, by Lemma˜2.8 we can take
$\Phi(P-P_{0})=L^{+}(\mathrm{cpt}_{\mathfrak{m}}(P-P_{0}))=\Phi(\mathrm{cpt}(\Sigma_{\mathfrak{m}})-\mathrm{cpt}_{\mathfrak{m}}(P_{0})),$
which shows the formula (2.13). ∎
Remark 2.28.
In case (b) of ˜2.27, the image of $Y(K)_{\Sigma_{\mathfrak{m}}}$ under the map (2.11) is contained in a translate of the group $\mathbb{Z}\cdot[\Sigma_{\mathfrak{m}}]\subseteq V_{D,\mathfrak{m}}$. This is a free subgroup of rank 1 unless $D$ consists of a single $K$-rational point, in which case $V_{D,\mathfrak{m}}$ itself is trivial. Indeed, if $\Sigma_{\mathfrak{m}}]=0$ in $V_{D,\mathfrak{m}}$, then $V_{D,\mathfrak{m}}=0$, so $\Sigma_{\mathfrak{m}}$ is the only closed point on $\tilde{\operatorname{\mathcal{D}}}$. As $\mathcal{X}$ is $D$-transversal at $\Sigma_{\mathfrak{m}}$, this shows that $D$ contains only a single point $Q$ and that $\tilde{\operatorname{\mathcal{D}}}=\operatorname{\mathcal{D}}=\mathcal{Q}$. The degree $[k(Q):K]$ equals the $\mathfrak{m}$-intersection number $i_{\mathfrak{m}}(\mathcal{X}_{\mathfrak{m}},\operatorname{\mathcal{D}})$, which must be equal to one as a consequence of the $D$-transversality condition and $\Sigma_{\mathfrak{m}}\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$.
Remark 2.29.
The $D$-transversality of $\mathcal{X}$ is important in the proof of ˜2.27. Otherwise, knowing that $P\in Y(K)$ reduces to $\Sigma_{\mathfrak{m}}$ on $\operatorname{\mathcal{D}}$ only constrains $\mathcal{P}.\tilde{\operatorname{\mathcal{D}}}$ to be an element of the subgroup generated by all points of $\tilde{\operatorname{\mathcal{D}}}$ lying over $\Sigma_{\mathfrak{m}}$, which could have rank larger than one.
Definition 2.30.
Let $P_{0}\in Y(K)$ be a base point. Let $(\mathcal{X},\operatorname{\mathcal{D}})$ be a regular $D$-transversal model of $(X,D)$ over $R$ and let $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}$ be a rational mod-$\mathfrak{m}$ reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$. We define the subset $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})\subseteq V_{D,\mathfrak{m}}$ as follows:
  1. (a)
    If $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{0}$ is a component, then $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$ is defined as the one-element subset of $V_{D,\mathfrak{m}}$ containing the right hand side of (2.12), i.e.
    $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})=\bigl\{-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(\Sigma_{\mathfrak{m}}-\mathrm{cpt}_{\mathfrak{m}}(P_{0})).\tilde{\operatorname{\mathcal{D}}}\bigr\}$. (2.14)
  2. (b)
    If $\Sigma_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}^{1}$ is a point in $(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$, then $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$ is defined as the set on the right hand side of (2.13), i.e.
    $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})=\mathbb{Z}\cdot[\Sigma_{\mathfrak{m}}]-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(\mathrm{cpt}(\Sigma_{\mathfrak{m}})-\mathrm{cpt}_{\mathfrak{m}}(P_{0})).\tilde{\operatorname{\mathcal{D}}}$. (2.15)
According to ˜2.27, we have
$(\sigma_{\mathfrak{m}}\circ\operatorname{AJ}_{P_{0}})(P)\in\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$
for all $P\in Y(K)$ of reduction type $\Sigma_{\mathfrak{m}}$.
Remark 2.31.
When $\Sigma_{\mathfrak{m}}$ is a component reduction type, it often happens that the singleton set $\mathfrak{S}(P_{0},\Sigma_{\mathfrak{m}})$ of (2.14) is contained in the $\mathfrak{S}(P_{0},\Sigma_{\mathfrak{m}}^{\prime})$ of (2.15) for a cuspidal reduction type $\Sigma_{\mathfrak{m}}^{\prime}$. Specifically, when $x\in\mathcal{C}_{\mathfrak{m}}^{1}$ is a point in $(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$ and $\mathrm{cpt}(x)\in\mathcal{C}_{\mathfrak{m}}^{0}$ is the component of the special fibre on which it lies, then $\mathfrak{S}_{\mathfrak{m}}(P_{0},\mathrm{cpt}(x))\subseteq\mathfrak{S}_{\mathfrak{m}}(P_{0},x)$. Moreover, if $\operatorname{\mathcal{D}}$ intersects the special fibre $\mathcal{X}_{\mathfrak{m}}$ in only one component, the set $\mathfrak{S}_{\mathfrak{m}}(P_{0},C)$ is in fact the same for each component $C\in\mathcal{C}_{\mathfrak{m}}^{0}$ by ˜2.16, so we have $\mathfrak{S}_{\mathfrak{m}}(P_{0},C)\subseteq\mathfrak{S}_{\mathfrak{m}}(P_{0},x)$ for all component reduction types $C\in\mathcal{C}_{\mathfrak{m}}^{0}$. These observations can be useful in applications of the Affine Chabauty method. They imply that components containing a point in $(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$ need not be considered separately as rational reduction types; and if $\operatorname{\mathcal{D}}$ intersects $\mathcal{X}_{\mathfrak{m}}$ in only one component and $(\mathcal{X}_{\mathfrak{m}}\cap\operatorname{\mathcal{D}})(k)\neq\emptyset$ (so that $\mathcal{C}_{\mathfrak{m}}^{1}\neq\emptyset$), no component reduction types need to be considered at all, as they are already covered by any cuspidal reduction type.
2.8. Comparison with the Kummer map
Assume that $K$ is a non-archimedean local field and $R$ is its valuation ring of residue characteristic $\ell$. Let $K^{s}/K$ be a separable closure and let $G_{K}=\operatorname{Gal}(K^{s}/K)$ be the absolute Galois group of $K$. Let $p$ be a prime. The short exact sequences of commutative group schemes
$0\longrightarrow J_{Y}[p^{n}]\longrightarrow J_{Y}\overset{\cdot p^{n}}{\longrightarrow}J_{Y}\longrightarrow 0$
induce the Kummer map
$\kappa\colon J_{Y}(K)\longrightarrow\mathrm{H}^{1}(G_{K},V_{p}J_{Y}),$ (2.16)
[…]
2.8.1. The case $\ell\neq p$
Here we show an “independence of pp” result for the Kummer map, namely that $V_{D,\mathfrak{m}}$ provides a $\mathbb{Q}$-form of the $\mathbb{Q}_{p}$-vector space $\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$, and that the Kummer map (2.16) factors through $V_{D,\mathfrak{m}}$ via the $D$-intersection map $\sigma_{\mathfrak{m}}$. Hence one may regard $\sigma_{\mathfrak{m}}$ as a “motivic” Kummer map of $J_{Y}$.
Theorem 2.32.
Assume $\ell\neq p$. Then there is a $\mathbb{Q}$-linear map $\varphi_{\mathbb{Q}}\colon V_{D,\mathfrak{m}}\to\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ inducing an isomorphism
$V_{D,\mathfrak{m}}\otimes_{\mathbb{Q}}\mathbb{Q}_{p}\cong\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$
and fitting into the commutative diagram
$J_{Y}(K)$$V_{D,\mathfrak{m}}$$\mathrm{H}^{1}(G_{K},V_{p}J_{Y}).$$\scriptstyle{\sigma_{\mathfrak{m}}}$$\scriptstyle{\kappa}$$\scriptstyle{\varphi_{\mathbb{Q}}}$ (2.17)
Proof.
[…]
The commutativity of (2.17) now follows from the functoriality of the Kummer maps and of $p$-adic completion together with the diagram
$T_{D}(K)$$J_{Y}(K)$$T_{D}(K)^{\wedge p}_{\mathbb{Q}_{p}}$$A_{D,\mathfrak{m}}$$\scriptstyle{\sigma_{\mathfrak{m}}}$$\scriptstyle{\beta}$
which is commutative by Lemma˜2.3. ∎
Now let $\mathcal{X}$ be a $D$-transversal model of $X$ over $R$. Let us compare the image of $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$ under $\varphi$ with the cohomological local Selmer set $\mathfrak{S}_{\mathfrak{m}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{m}})$ defined in [Bet23, §6.1]. If $\Sigma_{\mathfrak{m}}$ is an integral (resp. rational) reduction type, the latter is defined as the Zariski closure of the image of $\mathcal{Y}(R)_{\Sigma_{\mathfrak{m}}}$ (resp. $Y(K)_{\Sigma_{\mathfrak{m}}}$) under the composition $\kappa\circ\operatorname{AJ}_{P_{0}}\colon Y(K)\to J_{Y}(K)\to\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$, where the vector space $\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ is viewed as an affine $\mathbb{Q}_{p}$-scheme in the obvious way.
Corollary 2.33.
The set $\mathfrak{S}_{\mathfrak{m}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{m}})$ is equal to the Zariski closure of $\varphi_{\mathbb{Q}}(\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}}))$, so is either a singleton or an affine line. Moreover, inside $J_{Y}(K)$ we have
$\kappa^{-1}(\mathfrak{S}_{\mathfrak{m}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{m}}))=\sigma_{\mathfrak{m}}^{-1}(\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})).$ (2.18)
Proof.
The first and last assertion follow from the commutativity of (2.17) together with ˜2.13 and ˜2.27 once we know that $\sigma_{\mathfrak{m}}^{-1}(\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}}))$ is sufficiently large. If $\Sigma_{\mathfrak{m}}$ is non-cuspidal, it admits a smooth $k$-point, so by Hensel’s Lemma there exists an $R$-point $P\in\mathcal{Y}(R)$ reducing onto $\Sigma_{\mathfrak{m}}$, hence the singleton set $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$ (resp. $\mathfrak{S}_{\mathfrak{m}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{m}})$) contains $\sigma_{\mathfrak{m}}(\operatorname{AJ}_{P_{0}}(P))$ (resp. $\kappa(\operatorname{AJ}_{P_{0}}(P))$) as desired. Now assume that $\Sigma_{\mathfrak{m}}\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\mathcal{Q})(k)$ is a $k$-point on the horizontal extension of a cusp $Q$ defining a cuspidal reduction type. Let $t_{Q}$ be a local equation for $\mathcal{Q}$ near $\Sigma_{\mathfrak{m}}$, then $t_{Q}$ reduces to a uniformiser at $\Sigma_{\mathfrak{m}}$ by $D$-transversality, hence $t_{Q}$ defines a bijection between the residue disc of $\Sigma_{\mathfrak{m}}$ in $\mathcal{X}$ and $\mathfrak{m}$. Let $P$ be a $K$-rational point in this residue disc. Then
$i_{\Sigma_{\mathfrak{m}}}(\mathcal{P},\mathcal{Q})=\operatorname{ord}_{\Sigma_{\mathfrak{m}}}(t_{Q}|_{\mathcal{P}})=v_{K}(t_{Q}(P)),$
by [Lan88, Lemma III.2.2]. As $t_{Q}$ parametrises the residue disc, (2.13) shows that
$\sigma_{\mathfrak{m}}(\operatorname{AJ}_{P_{0}}(Y(K)_{\Sigma_{\mathfrak{m}}}))$ $=\mathbb{Z}_{>0}\cdot[\Sigma_{\mathfrak{m}}]-\mathcal{P}_{0}.\tilde{\operatorname{\mathcal{D}}}+\Phi(\mathrm{cpt}(\Sigma_{\mathfrak{m}})-\mathrm{cpt}_{\mathfrak{m}}(P_{0})).\tilde{\operatorname{\mathcal{D}}}\subseteq\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}}).$
We conclude using (2.15) and ˜2.32, which also shows that $\mathfrak{S}_{\mathfrak{m}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{m}})$ is an affine line. ∎
Remark 2.34.
˜2.33 reproves and sharpens [Bet23, Lemma 6.1.4] in the case $U=V_{p}J_{Y}$. Moreover, [BD20, Remark 1.1.3] prove an “independence of pp” result similar to ˜2.32 for the nonabelian Kummer map.
2.8.2. The case $\ell=p$
Now let $K/\mathbb{Q}_{p}$ be finite. For a $\mathbb{Q}_{p}$-linear representation $V$ of $G_{K}$, we write $\mathrm{H}^{1}_{f}(G_{K},V)\subseteq\mathrm{H}^{1}(G_{K},V)$ for the subspace of crystalline classes. We abbreviate $\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V)$ for $\mathrm{H}^{1}(G_{K},V)/\mathrm{H}^{1}_{f}(G_{K},V)$ and $\bar{\kappa}$ for the composition of $\kappa$ with $\mathrm{H}^{1}(G_{K},V)\to\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V)$.
Theorem 2.35.
There is a $\mathbb{Q}$-linear map $\varphi_{\mathbb{Q}}\colon V_{D,\mathfrak{m}}\to\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V_{p}J_{Y})$ making the diagram
$J_{Y}(K)$$V_{D,\mathfrak{m}}$$\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V_{p}J_{Y})$$\scriptstyle{\sigma_{\mathfrak{m}}}$$\scriptstyle{\bar{\kappa}}$$\scriptstyle{\varphi_{\mathbb{Q}}}$ (2.19)
commute.
Proof.
We again work with the integral structure $A_{D,\mathfrak{m}}$, construct a $\mathbb{Z}$-linear map $\varphi\colon A_{D,\mathfrak{m}}\to\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V_{p}J_{Y})$, and set $\varphi_{\mathbb{Q}}\coloneqq\varphi\otimes\mathbb{Q}$. We construct $\varphi$ as follows: Kummer theory for the extensions $k(Q)/K$ gives the Kummer isomorphism $\kappa\colon T_{D}(K)^{\wedge p}_{\mathbb{Q}_{p}}\xrightarrow{~\sim~}\mathrm{H}^{1}(G_{K},V_{p}T_{D})$ identifying crystalline classes with the $\kappa$-images of integral points. Thus $\bar{\kappa}\colon T_{D}(K)^{\wedge p}_{\mathbb{Q}_{p}}\to\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V_{p}T_{D})$ is given by taking valuations. Picking uniformisers for the fields $k(Q)$ gives a map $\beta\colon A_{D,\mathfrak{m}}\to T_{D}(K)^{\wedge p}_{\mathbb{Q}_{p}}$ such that $\bar{\kappa}\circ\beta\colon A_{D,\mathfrak{m}}\to T_{D}(K)^{\wedge p}_{\mathbb{Q}_{p}}$ is independent of the choice of uniformisers and an isomorphism after tensoring with $\mathbb{Q}_{p}$. Then define $\varphi\coloneqq\iota\circ\bar{\kappa}\circ\beta$, where $\iota$ is induced from $V_{p}T_{D}\to V_{p}J_{Y}$.
To show the commutativity of (2.19), we use that (2.1) gives the short exact sequence
$0\to T_{D}(K)^{\wedge p}_{\mathbb{Q}_{p}}\to J_{Y}(K)^{\wedge p}_{\mathbb{Q}_{p}}\to J(K)^{\wedge p}_{\mathbb{Q}_{p}}\to 0$ (2.20)
of finite dimensional $\mathbb{Q}_{p}$-vector spaces. For the Jacobian $J$, the Kummer map $\kappa\colon J(K)^{\wedge p}_{\mathbb{Q}_{p}}\to\mathrm{H}^{1}_{f}(G_{K},V_{p}J)$ is an isomorphism [BK90, p. 352] and hence $\bar{\kappa}\colon J(K)^{\wedge p}_{\mathbb{Q}_{p}}\to\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V_{p}J)$ is the zero map. Thus the image of any splitting $\chi\colon J(K)^{\wedge p}_{\mathbb{Q}_{p}}\to J_{Y}(K)^{\wedge p}_{\mathbb{Q}_{p}}$ of (2.20) is contained in the kernel of $\bar{\kappa}\colon J_{Y}(K)^{\wedge p}_{\mathbb{Q}_{p}}\to\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V_{p}J_{Y})$. As $\sigma_{\mathfrak{m}}\colon T_{D}(K)^{\wedge p}_{\mathbb{Q}_{p}}\to(A_{D,\mathfrak{m}})^{\wedge p}_{\mathbb{Q}_{p}}$ is surjective, we can choose a splitting $\chi$ of (2.20) in such a way that its image is contained in the kernel of $\sigma_{\mathfrak{m}}$. The commutativity of (2.19) now follows from the commutativity of
$J_{Y}(K)^{\wedge p}_{\mathbb{Q}_{p}}$$(A_{D,\mathfrak{m}})^{\wedge p}_{\mathbb{Q}_{p}}$$\mathrm{H}^{1}/\mathrm{H}^{1}_{f}(G_{K},V_{p}J_{Y}),$$\scriptstyle{\sigma_{\mathfrak{m}}}$$\scriptstyle{\bar{\kappa}}$$\scriptstyle{\varphi}$ (2.21)
[…]
where we used the functoriality of $\bar{\kappa}$ and that the image of $\chi$ is in the kernels of $\sigma_{\mathfrak{m}}$ and of $\bar{\kappa}$. ∎
The corollary below follows immediately from the commutativity of (2.19).
Corollary 2.36.
Inside $J_{Y}(K)$ we have
$\sigma_{\mathfrak{m}}^{-1}(\{0\})\subseteq\kappa^{-1}(\mathrm{H}^{1}_{f}(G_{K},V_{p}J_{Y})).$
3. Global theory: Selmer subsets of the generalised Jacobian
Let $K$ be a number field, let $X$ be a smooth projective curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. The ring of integers of $K$ is denoted by $\mathcal{O}_{K}$. We start by defining a $\mathbb{Q}$-vector space $V_{D}$ and a homomorphism $\sigma\colon J_{Y}(K)\to V_{D}$. In Section˜3.2 we define $S$-integral reduction types with respect to a regular model over $\mathcal{O}_{K}$ and use the map $\sigma$ to constrain the image of points of a fixed reduction type under an Abel–Jacobi map $Y\to J_{Y}$.
Definition 3.1.
Let $\pi\colon\tilde{\operatorname{\mathcal{D}}}\to\operatorname{Spec}(\mathcal{O}_{K})$ be the normalisation of $\operatorname{Spec}(\mathcal{O}_{K})$ in $D$. Concretely, if $D=\{Q_{1},\ldots,Q_{r}\}$ and $\mathcal{O}_{k(Q_{i})}$ denotes the ring of integers of the residue field $k(Q_{i})$, then
$\tilde{\operatorname{\mathcal{D}}}=\coprod_{i=1}^{r}\operatorname{Spec}(\mathcal{O}_{k(Q_{i})}).$
We also write $\tilde{\mathcal{Q}}_{i}=\operatorname{Spec}(\mathcal{O}_{k(Q_{i})})$.
3.1. The global $D$-intersection map
[…]
Definition 3.2.
Define the $\mathbb{Q}$-vector space
$V_{D}\coloneqq Z_{0}(\tilde{\operatorname{\mathcal{D}}})/\pi^{*}Z_{0}(\operatorname{Spec}(\mathcal{O}_{K}))\otimes_{\mathbb{Z}}\mathbb{Q}.$
The local $D$-intersection map of Section˜2.3 generalises to the global situation as follows. Choose a regular model $\mathcal{X}$ of $X$ over $\mathcal{O}_{K}$. With the same definitions as in Section˜2.2, we have intersection numbers $i_{x}(E,F)$ at closed points of $\mathcal{X}$ and we have an intersection cycle
$E.F=\sum_{x}i_{x}(E,F)[x]$
for any two ($\mathbb{Q}$-)divisors on $\mathcal{X}$ without common component in their support. We also have $\mathfrak{q}$-intersection numbers $i_{\mathfrak{q}}(E,F)$ for all primes $\mathfrak{q}$ of $\mathcal{O}_{K}$, defined whenever $E$ and $F$ have no common horizontal component in their support. Via the valuative criterion of properness, the inclusion $D\hookrightarrow\mathcal{X}$ extends to a morphism of $\mathcal{O}_{K}$-schemes
$\nu\colon\tilde{\operatorname{\mathcal{D}}}\to\mathcal{X},$
[…]
Lemma 3.3.
For any degree-0 divisor $F$ on $Y$ with horizontal extension $\mathcal{F}$ on $\mathcal{X}$ there exists a vertical $\mathbb{Q}$-divisor $\Phi(F)$ on $\mathcal{X}$ such that for all primes $\mathfrak{q}$ of $\mathcal{O}_{K}$, the divisor $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$ has $\mathfrak{q}$-intersection number zero with all vertical divisors over $\mathfrak{q}$. The divisor $\Phi(F)$ is unique up to adding $\mathbb{Q}$-divisors which are pulled back along $\mathcal{X}\to\operatorname{Spec}(\mathcal{O}_{K})$.
Proof.
By [Lan88, Theorem III.3.6], there exists, for each prime $\mathfrak{q}$ of $\mathcal{O}_{K}$, a vertical $\mathbb{Q}$-divisor $\Phi_{\mathfrak{q}}(F)$ on $\mathcal{X}$ over the prime $\mathfrak{q}$ such that $\Psi_{\mathfrak{q}}(F)\coloneqq\mathcal{F}+\Phi_{\mathfrak{q}}(F)$ has $\mathfrak{q}$-intersection number zero with every vertical divisor on $\mathcal{X}$ over $\mathfrak{q}$. For primes $\mathfrak{q}$ where $\mathcal{X}$ has good reduction, we can take $\Phi_{\mathfrak{q}}(F)=0$. Then $\Phi(F)\coloneqq\sum_{\mathfrak{q}}\Phi_{\mathfrak{q}}(F)$ has the required properties. The uniqueness statement follows from the fact that $\Phi_{\mathfrak{q}}(F)$ is unique up to adding $\mathbb{Q}$-multiples of the complete fibre $\mathcal{X}_{\mathfrak{q}}$. ∎
We first define $\sigma\colon J_{Y}(K)\to V_{D}$ on the level of divisors. Given a degree-0 divisor $F$ on $Y$, choose $\Phi(F)$ as in Lemma˜3.3 and set $\Psi(F)\coloneqq\mathcal{F}+\Phi(F)$. Then define
$\sigma(F)\coloneqq\Psi(F).\tilde{\operatorname{\mathcal{D}}}$, (3.1)
which is well-defined modulo $\pi^{*}Z_{0}(\operatorname{Spec}(\mathcal{O}_{K}))$ and hence as an element of $V_{D}$ by the uniqueness statement in Lemma˜3.3. The global variant of Lemma˜2.3 reads as follows:
Lemma 3.4.
Let $F=\mathrm{div}(f)$ be a principal divisor with support in $Y$. Then we have
$\sigma(F)=\sum_{Q\in|D|}\;\sum_{x\in|\tilde{\mathcal{Q}}|}\operatorname{ord}_{x}(f(Q))[x],$
where in the inner sum $x$ runs over the closed points of $\tilde{\mathcal{Q}}=\operatorname{Spec}(\mathcal{O}_{k(Q)})$.
Proof.
We can choose $\Psi(F)=\mathrm{div}_{\mathcal{X}}(f)$ and proceed as in the proof of Lemma˜2.3. ∎
As in the local case, it follows from Lemma˜3.4 that the map (3.1) is well-defined on $D$-equivalence classes of divisors, i.e. it descends to $J_{Y}(K)$.
Definition 3.5.
We call the homomorphism $\sigma\colon J_{Y}(K)\to V_{D}$, $F]\mapsto\Psi(F).\tilde{\operatorname{\mathcal{D}}}$ the global $D$-intersection map.
˜2.6 easily carries over to the global situation, so the global $D$-intersection map is also independent of the choice of regular model $\mathcal{X}$. The global map $\sigma$ can be viewed as the direct sum of local maps $\sigma_{\mathfrak{q}}$ as follows. We have a natural isomorphism $Z_{0}(\tilde{\operatorname{\mathcal{D}}})\cong\bigoplus_{\mathfrak{q}}Z_{0}(\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{q}})$, with $\mathfrak{q}$ running over all primes of $\mathcal{O}_{K}$ and $\operatorname{\mathcal{D}}_{\mathfrak{q}}$ denoting the fibre of $\operatorname{\mathcal{D}}$ over $\mathfrak{q}$. With $V_{D,\mathfrak{q}}\coloneqq Z_{0}(\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{q}})/[\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{q}}]\otimes_{\mathbb{Z}}\mathbb{Q}$ this induces an isomorphism
$V_{D}\cong\bigoplus_{\mathfrak{q}}V_{D,\mathfrak{q}}.$ (3.2)
For each prime $\mathfrak{q}$ of $\mathcal{O}_{K}$, the theory of Section˜2 applied to the (uncompleted) discrete valuation ring $\mathcal{O}_{K,\mathfrak{q}}$ yields a local $D$-intersection map $\sigma_{\mathfrak{q}}\colon J_{Y}(K)\to V_{D,\mathfrak{q}}$.
Proposition 3.6.
Under the identification (3.2), the global $D$-intersection map $\sigma\colon J_{Y}(K)\to V_{D}$ equals the sum of the local maps $\sigma_{\mathfrak{q}}\colon J_{Y}(K)\to V_{D,\mathfrak{q}}$ over all primes $\mathfrak{q}$ of $\mathcal{O}_{K}$.
Proof.
If $\mathcal{X}$ is a regular model of $X$ over $\mathcal{O}_{K}$, then for each prime $\mathfrak{q}$ of $\mathcal{O}_{K}$, the base change $\mathcal{X}_{\mathcal{O}_{K,\mathfrak{q}}}=\mathcal{X}\otimes_{\mathcal{O}_{K}}\mathcal{O}_{K,\mathfrak{q}}$ is a regular model over $\mathcal{O}_{K,\mathfrak{q}}$. The maps $\nu_{\mathfrak{q}}\colon\tilde{\operatorname{\mathcal{D}}}_{\mathcal{O}_{K,\mathfrak{q}}}\to\mathcal{X}_{\mathcal{O}_{K,\mathfrak{q}}}$ used to define the generalised intersection numbers on $\mathcal{X}_{\mathcal{O}_{K,\mathfrak{q}}}$ are obtained by base change from $\nu\colon\tilde{\operatorname{\mathcal{D}}}\to\mathcal{X}$. For a degree-0 divisor $F$ on $Y$ and $\Psi(F)=\mathcal{F}+\sum_{\mathfrak{q}}\Phi_{\mathfrak{q}}$ as in Lemma˜3.3, the part of $\sigma(F)=\Psi(F).\tilde{\operatorname{\mathcal{D}}}$ belonging to the fibre over $\mathfrak{q}$ is $(\mathcal{F}+\Phi_{\mathfrak{q}}(F)).\tilde{\operatorname{\mathcal{D}}}_{\mathcal{O}_{K,\mathfrak{q}}}$, which is precisely $\sigma_{\mathfrak{q}}(F)$. ∎
Proposition 3.7.
The kernel of $\sigma\colon J_{Y}(K)\to V_{D}$ is finitely generated of rank
$\operatorname{\mathrm{rk}}\ker(\sigma)=n_{1}(D)+n_{2}(D)-\#|D|-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}+r,$
where $n_{1}(D)=\#D(\mathbb{R})$ is the number of real cusps, $n_{2}(D)$ is the number of conjugate pairs of complex cusps, and $r=\operatorname{\mathrm{rk}}J(K)$ is the Mordell–Weil rank of the Jacobian of  $X$.
Proof.
Recall from Section˜2.1 that the generalised Jacobian fits into a short exact sequence
$0\to T_{D}\to J_{Y}\to J\to 0,$
which remains exact on the level of $K$-points. The $K$-points of the torus $T_{D}$ are $T_{D}(K)=\mathcal{O}(D)^{\times}/K^{\times}$. An element of $\mathcal{O}(D)^{\times}=\prod_{Q\in|D|}k(Q)^{\times}$ defines a zero-cycle on each $\tilde{\mathcal{Q}}$, hence a zero-cycle on $\tilde{\operatorname{\mathcal{D}}}$. Elements of $K^{\times}$ give rise to zero-cycles which are pulled back from $\operatorname{Spec}(\mathcal{O}_{K})$. So we get a well-defined divisor map $\mathrm{div}\colon T_{D}(K)\to V_{D}$. This gives rise to the following diagram with exact rows:
[…]
The left square is commutative. This follows from Lemma˜3.4 and the fact that the map $T_{D}(K)\to J_{Y}(K)$ sends a tuple $(a_{Q})_{Q}\in\prod k(Q)^{\times}$ to the class of the principal divisor $\mathrm{div}(f)$ defined by any rational function $f$ on $X$ that interpolates the values $f(Q)=a_{Q}$ for $Q\in|D|$. The snake lemma yields an exact sequence
$0\to\ker(\mathrm{div})\to\ker(\sigma)\to J(K)\to\operatorname{coker}(\mathrm{div})\to\operatorname{coker}(\sigma)\to 0.$ (3.3)
We claim that $\operatorname{coker}(\mathrm{div})$ is torsion. Let $c$ be a zero-cycle on $\tilde{\operatorname{\mathcal{D}}}$ with $\mathbb{Q}$-coefficients, representing an element of $V_{D}$. Let $m$ be a positive integer such that $mc$ has $\mathbb{Z}$-coefficients. Write $mc=\sum_{Q\in|D|}c_{Q}^{\prime}$ with $c_{Q}^{\prime}$ a zero-cycle on $\tilde{\mathcal{Q}}=\operatorname{Spec}\mathcal{O}_{k(Q)}$. Let $h$ be the least common multiple of the class numbers $\mathrm{Cl}(k(Q))$ for $Q\in|D|$, so that $hc_{Q}^{\prime}$ is a principal divisor on $\tilde{\mathcal{Q}}$, say $\mathrm{div}(a_{Q})$ with $a_{Q}\in k(Q)^{\times}$. Then mhcmhc is the image of $(a_{Q})_{Q}$ under $\mathrm{div}\colon T_{D}(K)\to V_{D}$, hence $c$ is killed by $mh$ in $\operatorname{coker}(\mathrm{div})$. This shows that the cokernel is torsion. Since $J(K)$ is finitely generated, its image in $\operatorname{coker}(\mathrm{div})$ is finite, so the exact sequence (3.3) implies
$\operatorname{\mathrm{rk}}\ker(\sigma)=\operatorname{\mathrm{rk}}\ker(\mathrm{div})+\operatorname{\mathrm{rk}}J(K).$ (3.4)
We are reduced to determining the rank of $\ker(\mathrm{div})$. Consider the following commutative diagram with exact rows
[…]
The snake lemma yields an exact sequence
$0\to\mathcal{O}_{K}^{\times}\to\mathcal{O}(\tilde{\operatorname{\mathcal{D}}})^{\times}\to\ker(\mathrm{div})\to\mathrm{Cl}(K).$
Since the class group $\mathrm{Cl}(K)$ is finite, we get
$\operatorname{\mathrm{rk}}\ker(\mathrm{div})=\operatorname{\mathrm{rk}}\mathcal{O}(\tilde{\operatorname{\mathcal{D}}})^{\times}-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}.$ (3.5)
[…]
$\operatorname{\mathrm{rk}}\mathcal{O}(\tilde{\operatorname{\mathcal{D}}})^{\times}=\sum_{Q\in|D|}\operatorname{\mathrm{rk}}\mathcal{O}_{k(Q)}^{\times}=\sum_{Q\in|D|}(n_{1}(Q)+n_{2}(Q)-1)=n_{1}(D)+n_{2}(D)-\#|D|.$ (3.6)
The claimed rank of $\ker(\sigma)$ is obtained by combining (3.4), (3.5), (3.6). ∎
3.2. $S$-integral reduction types
Let $S$ be a finite set of primes of $\mathcal{O}_{K}$. Let $\mathcal{X}$ be a regular model of $X$ over the ring of integers $\mathcal{O}_{K}$, let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. We assume that $\mathcal{X}$ is $D$-transversal over primes in $S$, i.e. the base change $\mathcal{X}_{\mathcal{O}_{K,\mathfrak{q}}}$ to the local ring $\mathcal{O}_{K,\mathfrak{q}}$ is $D$-transversal for all $\mathfrak{q}\in S$ (see ˜2.17). The set of $S$-integral points $\mathcal{Y}(\mathcal{O}_{K,S})$ can be partitioned according to their mod-$\mathfrak{q}$ reductions for all primes $\mathfrak{q}$ as follows.
Definition 3.8.
An $S$-integral reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$ is a collection $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ consisting of
  • for each prime $\mathfrak{q}\not\in S$: an integral mod-$\mathfrak{q}$ reduction type $\Sigma_{\mathfrak{q}}$ for $(\mathcal{X},\operatorname{\mathcal{D}})$, see ˜2.11;
  • for each prime $\mathfrak{q}\in S$: a rational mod-$\mathfrak{q}$ reduction type $\Sigma_{\mathfrak{q}}$ for $(\mathcal{X},\operatorname{\mathcal{D}})$, see ˜2.25.
An $S$-integral point $P\in\mathcal{Y}(\mathcal{O}_{K,S})$ has reduction type $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ if for all primes $\mathfrak{q}$ of $\mathcal{O}_{K}$, the image of $P$ under the mod-$\mathfrak{q}$ reduction map ${\mathrm{red}}_{\mathfrak{q}}\colon\mathcal{Y}(\mathcal{O}_{K,S})\hookrightarrow\mathcal{X}(\mathcal{O}_{K,S})=\mathcal{X}(\mathcal{O}_{K})\to\mathcal{X}(\operatorname{\mathbb{F}}_{\mathfrak{q}})$ lies on the component resp. is equal to the point $\Sigma_{\mathfrak{q}}$. The set of all $S$-integral points of reduction type $\Sigma$ is denoted by $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$.
This defines a partition
$\mathcal{Y}(\mathcal{O}_{K,S})=\coprod_{\Sigma}\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$
with $\Sigma$ running over all $S$-integral reduction types for $(\mathcal{X},\operatorname{\mathcal{D}})$. For a prime $\mathfrak{q}$ where $\mathcal{X}$ has good reduction and which is not contained in $S$, there is only one possible choice for $\Sigma_{\mathfrak{q}}$, namely the unique component of the mod-$\mathfrak{q}$ fibre $\mathcal{X}_{\mathfrak{q}}\smallsetminus\operatorname{\mathcal{D}}$. In particular, there are only finitely many $S$-integral reduction types for $(\mathcal{X},\operatorname{\mathcal{D}})$.
3.3. Selmer sets
Let $\mathcal{X}/\mathcal{O}_{K}$ be as before and fix an $S$-integral reduction type $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ for $(\mathcal{X},\operatorname{\mathcal{D}})$. We also fix a base point $P_{0}\in Y(K)$.
Recall from §3.1 the $\mathbb{Q}$-vector space $V_{D}=Z_{0}(\tilde{\operatorname{\mathcal{D}}})/\pi^{*}Z_{0}(\operatorname{Spec}(\mathcal{O}_{K}))\otimes_{\mathbb{Z}}\mathbb{Q}$ and its direct sum decomposition into the local vector spaces $V_{D,\mathfrak{q}}=Z_{0}(\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{q}})/[\tilde{\operatorname{\mathcal{D}}}_{\mathfrak{q}}]\otimes_{\mathbb{Z}}\mathbb{Q}$ with $\mathfrak{q}$ running over all primes $\mathfrak{q}$ of $\mathcal{O}_{K}$:
$V_{D}\cong\bigoplus_{\mathfrak{q}}V_{D,\mathfrak{q}}.$
For all primes $\mathfrak{q}$ we have constructed a subset $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})\subseteq V_{D,\mathfrak{q}}$, see ˜2.14 (if $\mathfrak{q}\not\in S$) and ˜2.30 (if $\mathfrak{q}\in S$). They provide the local conditions to define a global subset $\mathfrak{S}(P_{0},\Sigma)$ of $V_{D}$.
Definition 3.9.
Define the subset $\mathfrak{S}(P_{0},\Sigma)$ of $V_{D}$ by
$\mathfrak{S}(P_{0},\Sigma)\coloneqq\prod_{\mathfrak{q}}\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}}).$
In other words, an element $c=(c_{\mathfrak{q}})_{\mathfrak{q}}$ of $V_{D}$ belongs to $\mathfrak{S}(P_{0},\Sigma)$ if and only $c_{\mathfrak{q}}\in\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})$ for all primes $\mathfrak{q}$.
Note that $\prod_{\mathfrak{q}}\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})$ is indeed contained in the direct sum of the vector spaces $V_{D,\mathfrak{q}}$ since it follows from ˜2.15 that $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})=\{0\}$ for all primes $\mathfrak{q}\not\in S$ such that $\mathcal{X}$ has good reduction at $\mathfrak{q}$ and $P_{0}$ is integral at $\mathfrak{q}$.
Definition 3.10.
Denote by $C(\Sigma)$ the set of primes $\mathfrak{q}\in S$ for which $\Sigma_{\mathfrak{q}}$ is cuspidal, i.e. an element of $\mathcal{C}_{\mathfrak{q}}^{1}=(\mathcal{X}_{\mathfrak{q}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\mathfrak{q}})$. The cuspidal part $\Sigma^{{\mathrm{csp}}}$ of $\Sigma$ is its image under the projection to $\prod_{\mathfrak{q}\in C(\Sigma)}\mathcal{C}_{\mathfrak{q}}^{1}$.
Proposition 3.11.
Assume that $D$ does not consist of a single $K$-rational point. Then $\mathfrak{S}(P_{0},\Sigma)$ is a translate of a subgroup $U(\Sigma^{{\mathrm{csp}}})$ of $V_{D}$ which is free of rank $C(\Sigma)$.
Proof.
The assumption on $D$ ensures that the vector spaces $V_{D}$ and $V_{D,\mathfrak{q}}$ are non-trivial. For $\mathfrak{q}\in C(\Sigma)$, the set $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})$ is a translate of the rank 1 subgroup $\mathbb{Z}\cdot[\Sigma_{\mathfrak{q}}]\subseteq V_{D,\mathfrak{q}}$, see ˜2.28; for all other $\mathfrak{q}$ it is a singleton. ∎
Definition 3.12.
Define the Selmer set associated to the reduction type $\Sigma$ and base point $P_{0}$ as the preimage of $\mathfrak{S}(P_{0},\Sigma)$ in $J_{Y}(K)$ under the $D$-intersection map:
$\operatorname{Sel}(P_{0},\Sigma)\coloneqq\sigma^{-1}(\mathfrak{S}(P_{0},\Sigma))\subseteq J_{Y}(K).$
Proposition 3.13.
The Abel–Jacobi map $\operatorname{AJ}_{P_{0}}\colon Y(K)\to J_{Y}(K)$ maps all points of reduction type $\Sigma$ into $\operatorname{Sel}(P_{0},\Sigma)$:
$\operatorname{AJ}_{P_{0}}(\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma})\subseteq\operatorname{Sel}(P_{0},\Sigma).$
Proof.
If $P$ has reduction type $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ then, by definition, its mod-$\mathfrak{q}$ reduction type is $\Sigma_{\mathfrak{q}}$, for all primes $\mathfrak{q}$ of $\mathcal{O}_{K}$. By the local theory, specifically ˜2.13 and ˜2.27, this implies $\sigma_{\mathfrak{q}}(\operatorname{AJ}_{P_{0}}(P))\in\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})$ for all $\mathfrak{q}$. Since $\sigma_{\mathfrak{q}}$ equals the $\mathfrak{q}$-component of the global map $\sigma$ (see ˜3.6), the claim follows. ∎
Proposition 3.14.
Assume that $D$ does not consist of a single $K$-rational point. The Selmer set $\operatorname{Sel}(P_{0},\Sigma)$ is either empty or a translate of a subgroup of $J_{Y}(K)$ of rank $n_{1}(D)+n_{2}(D)-\#|D|-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}+r+\#C(\Sigma)$.
Proof.
By ˜3.11, the set $\mathfrak{S}(P_{0},\Sigma)$ is a translate of a subgroup $U$ of $V_{D}$ which has rank $C(\Sigma)$. If $\operatorname{Sel}(P_{0},\Sigma)$ is not empty, then it is a translate of the subgroup $\sigma^{-1}(U)$ of $J_{Y}(K)$. It remains to determine the rank of $\sigma^{-1}(U)$. Let $\sigma_{U}:\sigma^{-1}(U)\to U$ be the restriction of $\sigma$ and consider the short exact sequence
$0\longrightarrow\ker(\sigma)\longrightarrow\sigma^{-1}(U)\overset{\sigma_{U}}{\longrightarrow}U\longrightarrow\operatorname{coker}(\sigma_{U})\longrightarrow 0.$ (3.7)
It was shown in the proof of ˜3.7 that the cokernel of $\mathrm{div}\colon T_{D}(K)\to V_{D}$ and thus also the cokernel of $\sigma$ are torsion. Since $\operatorname{coker}(\sigma_{U})$ embeds into $\operatorname{coker}(\sigma)$, the cokernel of $\sigma_{U}$ is torsion as well. As a quotient of $U$ it is also finitely generated, hence finite. Now the exact sequence (3.7) shows
$\operatorname{\mathrm{rk}}\sigma^{-1}(U)=\operatorname{\mathrm{rk}}\ker(\sigma)+\operatorname{\mathrm{rk}}U$
and the claim follows from $\operatorname{\mathrm{rk}}U=\#C(\Sigma)$ and the computation of the rank of $\ker(\sigma)$ in ˜3.7. ∎
Corollary 3.15.
Let $\mathcal{X}/\mathcal{O}_{K,S}$ be a regular model of $X$ over the ring of $S$-integers. Fix a base point $P_{0}\in Y(K)$. The image of the set of $S$-integral points $\mathcal{Y}(\mathcal{O}_{K,S})$ under the Abel–Jacobi embedding $\operatorname{AJ}_{P_{0}}\colon Y(K)\to J_{Y}(K)$ is contained in a finite union of translates of subgroups of rank at most $n_{1}(D)+n_{2}(D)-\#|D|-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}+r+\#S$.
Proof.
We can assume that the set of cusps $D$ does not consist only of a single $K$-rational point, for otherwise the generalised Jacobian of $Y$ coincides with the Jacobian of the compactification $X$, whose group of $K$-points has rank $r$ by definition. Let $\mathcal{X}$ be a regular model of $X$ and extend it to a regular model over $\operatorname{Spec}(\mathcal{O}_{K})$ that is $D$-transversal over primes in $S$, compare Section˜2.6. The set $\mathcal{Y}(\mathcal{O}_{K,S})$ is partitioned into finitely many subsets $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$ by reduction types (˜3.8), and each non-empty $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$ is mapped under the Abel–Jacobi map to a translate of a subgroup of rank $n_{1}(D)+n_{2}(D)-\#|D|-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}+r+\#C(\Sigma)$ by ˜3.14. Here, $C(\Sigma)$ is the set of primes $\mathfrak{q}\in S$ at which the local reduction type $\Sigma_{\mathfrak{q}}$ is a point on the closure of a cusp. In particular, $C(\Sigma)$ is at most #S\#S. ∎
Remark 3.16.
We compare the Selmer set $\operatorname{Sel}(P_{0},\Sigma)\subseteq J_{Y}(K)$ with the cohomological Selmer set $\operatorname{Sel}_{\Sigma,U_{1}}\subseteq\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ of [Bet23, §6.2], where $U_{1}=V_{p}J_{Y}$ is the abelianisation of the $\mathbb{Q}_{p}$-prounipotent fundamental group of $Y$. Assume that all primes $\mathfrak{q}|p$ are primes of good reduction for $(\mathcal{X},\operatorname{\mathcal{D}})$, that $P_{0}$ is integral at those primes, and that they are not contained in $S$. The set $\operatorname{Sel}_{\Sigma,U_{1}}$ is defined as the set of those global cohomology classes whose restriction at primes $\mathfrak{q}|p$ is crystalline, and whose restriction lies in $\mathfrak{S}^{\mathrm{coh}}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})\subseteq\mathrm{H}^{1}(G_{K_{\mathfrak{q}}},V_{p}J_{Y})$ for all primes $\mathfrak{q}\nmid p$; see §2.8 for the definition of $\mathfrak{S}_{\mathfrak{q}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{q}})$. ([Bet23, §6] also assumes that the base point $P_{0}$ is $S$-integral and that $K=\mathbb{Q}$ but the definition makes sense more generally.) Our assumptions ensure that there is only one reduction type $\Sigma_{\mathfrak{q}}$ at primes $\mathfrak{q}\mid p$ and $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})=\{0\}$. Let $\kappa\colon J_{Y}(K)\to\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ denote the global Kummer map, which is defined in the same way as in (2.16) and compatible with the local Kummer maps. Now ˜2.33 and ˜2.36 show
$\operatorname{Sel}(P_{0},\Sigma)\subseteq\kappa^{-1}(\operatorname{Sel}_{\Sigma,U_{1}}).$ (3.8)
The Selmer set $\operatorname{Sel}(P_{0},\Sigma)$ can thus be viewed as a geometric incarnation of the cohomological Selmer set. In this sense our method refines the abelian affine case of [Bet23] that was also considered in [LLM23, §4]. In particular, our method is at least as strong as Kim’s non-abelian Chabauty method in depth 1 in its refined version by Betts–Dogra [BD20].
3.4. Reduction types for more general models
Let $S$ be a finite set of primes of $\mathcal{O}_{K}$. Let $Y=X\smallsetminus D$ as before and let $\mathcal{X}$ be a model of $X$ over $\mathcal{O}_{K}$ [Liu06, Definition 10.1.1] which is not necessarily regular and not necessarily $D$-transversal over primes in $S$. Let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. We can still partition the set of $S$-integral points $\mathcal{Y}(\mathcal{O}_{K,S})$ into subsets whose images under an Abel–Jacobi map $Y(K)\to J_{Y}(K)$ are strongly constrained, but we have to use reduction types on a suitably desingularised model. Let $\pi\colon\mathcal{X}^{\prime}\to\mathcal{X}$ be a morphism of models over $\mathcal{O}_{K}$ where $\mathcal{X}^{\prime}$ is regular and $D$-transversal over primes in $S$. Such a desingularisation can be obtained as follows: by repeatedly blowing up in the singular locus and normalising, one obtains a regular model [Liu06, Cor. 8.3.51], and then one can use ˜2.19. Let $\operatorname{\mathcal{D}}^{\prime}$ be the closure of $D$ in $\mathcal{X}^{\prime}$ and set $\mathcal{Y}^{\prime}=\mathcal{X}^{\prime}\smallsetminus\operatorname{\mathcal{D}}^{\prime}$. We have the inclusion $\mathcal{Y}(\mathcal{O}_{K,S})\subseteq\mathcal{Y}^{\prime}(\mathcal{O}_{K,S})$, which can be strict, for example if $\pi$ contracts a vertical component onto a point lying on $\operatorname{\mathcal{D}}$. But by allowing only certain reduction types, we can find $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\mathcal{Y}^{\prime}(\mathcal{O}_{K,S})$ and thereby reduce the computation of $S$-integral points on $\mathcal{Y}$ to the case of a model which is regular and $D$-transversal over primes in $S$.
Definition 3.17.
An $S$-integral reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$ (specified on $\mathcal{X}^{\prime}$) is a collection $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ where $\Sigma_{\mathfrak{q}}$ is:
  • for primes $\mathfrak{q}\not\in S$: an irreducible (equivalently: connected) component of ${\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\smallsetminus\pi^{-1}(\operatorname{\mathcal{D}})$ containing an $\operatorname{\mathbb{F}}_{\mathfrak{q}}$-point;
  • for primes $\mathfrak{q}\in S$: either a component of ${\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}^{\prime}$ containing an $\operatorname{\mathbb{F}}_{\mathfrak{q}}$-point, or a point in $({\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(\operatorname{\mathbb{F}}_{\mathfrak{q}})$.
For a prime $\mathfrak{q}$, the composition
$Y(K)\overset{{\mathrm{red}}_{\mathfrak{q}}}{\longrightarrow}{\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}(\operatorname{\mathbb{F}}_{\mathfrak{q}})\overset{\pi}{\longrightarrow}\mathcal{X}_{\mathfrak{q}}(\operatorname{\mathbb{F}}_{\mathfrak{q}})$
of the mod-$\mathfrak{q}$ reduction map for the model $\mathcal{X}^{\prime}$ with the morphism $\pi\colon\mathcal{X}^{\prime}\to\mathcal{X}$ equals the mod-$\mathfrak{q}$ reduction map for $\mathcal{X}$, taking values in possibly singular $\operatorname{\mathbb{F}}_{\mathfrak{q}}$-points of $\mathcal{X}_{\mathfrak{q}}$. It is clear from the definition that the $S$-integral points of $\mathcal{Y}$ are exactly those $K$-points which have an $S$-integral reduction type for $(\mathcal{X},\operatorname{\mathcal{D}})$ as defined above. Thus, computing $\mathcal{Y}(\mathcal{O}_{K,S})$ is equivalent to computing $\mathcal{Y}^{\prime}(\mathcal{O}_{K,S})_{\Sigma}$ for all reduction types $\Sigma$ which are $S$-integral for $(\mathcal{X},\operatorname{\mathcal{D}})$.
4. $p$-adic integrals of logarithmic differentials
Let $K$ be a complete subfield of $\mathbb{C}_{p}$ (such as $\mathbb{Q}_{p}$ or a finite extension thereof), let $X$ be a smooth projective curve of genus $g$ over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. Denote by $n\coloneqq D(\overline{K})$ the number of geometric cusps. A logarithmic differential on $(X,D)$ is a meromorphic differential on $X$ which is everywhere regular except for possibly simple poles at the points of $D$. The $K$-vector space of logarithmic differentials on $(X,D)$ is $\mathrm{H}^{0}(X,\Omega^{1}(D))$. By the Riemann–Roch theorem, its dimension is given by
$\dim_{K}\mathrm{H}^{0}(X,\Omega^{1}(D))=g+n-1.$ (4.1)
4.1. The abelian integral
We recall the construction of $p$-adic integrals of log differentials. There are different possible approaches; for our purposes it is most convenient to work with the abelian integral constructed by Zarhin [Zar96] and Colmez [Col98]. With this approach it is not necessary to assume that $X$ has good reduction.
The abelian integral is first constructed on the generalised Jacobian $J_{Y}$ using the logarithm map of the $p$-adic Lie group $J_{Y}(K)$, and then pulled back to $Y(K)$. The construction by Zarhin [Zar96] requires the choice of a branch of the $p$-adic logarithm $\log\colon K^{\times}\to K$. This choice is uniquely determined by specifying $\log(p)\in K$. The construction by Colmez [Col98] avoids such a choice by treating $\log(p)$ as a formal symbol, so that the abelian integrals have values in $K_{\mathrm{st}}\coloneqq K[\log(p)]$. This specialises to Zarhin’s definition upon choosing a value for $\log(p)\in K$. For simplicity, we choose to work with the Iwasawa logarithm, defined by $\log(p)=0$. Once this choice is made, the logarithm map of the Lie group $J_{Y}(K)$, which a priori is only defined on an open neighbourhood of the identity, can be extended to all of $J_{Y}(K)$, resulting in a homomorphism
$\log_{J_{Y}}\colon J_{Y}(K)\to\operatorname{Lie}(J_{Y}).$
The Lie algebra of $J_{Y}$ is the dual of the space of invariant differential forms on $J_{Y}$. Assume that $Y(K)\neq\emptyset$. Pullback along an Abel–Jacobi map $Y\to J_{Y}$ defines an isomorphism between invariant differentials on $J_{Y}$ and log differentials on $(X,D)$. This isomorphism does not depend on the base point of the Abel–Jacobi map. So the logarithm of $J_{Y}$ can be viewed as a map
$\log_{J_{Y}}\colon J_{Y}(K)\to\mathrm{H}^{0}(X,\Omega^{1}(D))^{\vee}.$ (4.2)
This allows us to define, for any degree zero divisor $F$ on $\operatorname{Div}^{0}(Y)$, representing an element of $J_{Y}(K)$, and for any log differential $\omega$ on $(X,D)$ the integral
$\int_{F}\omega\coloneqq\log_{J_{Y}}(F)(\omega)\in K.$
Pulling this back to $Y$, we can now define
$\int_{P}^{Q}\omega\coloneqq\int_{[Q]-[P]}\omega$
for any two points $P,Q\in Y(K)$ and $\omega\in\mathrm{H}^{0}(X,\Omega^{1}(D))$.
Example 4.1.
On $Y=\mathbb{G}_{m}=\mathbb{P}^{1}\smallsetminus\{0,\infty\}$ we have
$\int_{a}^{b}\frac{\mathrm{d}z}{z}=\log\left(\frac{b}{a}\right)$
for all $a,b\in K^{\times}$, where $\log\colon K^{\times}\to K$ is the chosen branch of the $p$-adic logarithm.
4.2. Tiny integrals
Assume that $K$ is a finite extension of $\mathbb{Q}_{p}$. Let $R$ be the valuation ring of $K$ with maximal ideal $\mathfrak{m}$ and finite residue field $k$. Let $\mathcal{X}$ be a regular model of $X$ over $R$, let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. We have the mod-$\mathfrak{m}$ reduction map
${\mathrm{red}}_{\mathfrak{m}}\colon\mathcal{Y}(R)\to\mathcal{Y}_{\mathfrak{m}}^{\mathrm{sm}}(k)$
taking values in the smooth $k$-points of the special fibre, see ˜2.10. The fibres of this map are called residue discs. A $p$-adic integral whose endpoints lie in the same residue disc is called a tiny integral. They can be computed by formally integrating a power series, as we now explain.
[…]
$t\colon U(\tilde{P})\xrightarrow{\sim}\mathfrak{m}.$
[…]
$\int_{P}^{Q}\omega=F(t(Q))-F(t(P)).$ (4.3)
In particular, if $\omega\neq 0$ and if $P$ is fixed, then the function $Q\mapsto\int_{P}^{Q}\omega$ has only finitely many zeros inside the disc U(P~)U($\tilde${P}). Indeed, by (4.3), the function $Q\mapsto\int_{P}^{Q}\omega$ is given by evaluating the power series $F$ (plus a constant) at elements of $\mathfrak{m}$, and looking at the Newton polygon gives the result. For a more precise version, see Lemma˜6.3.
5. Construction of Chabauty functions
We now prove ˜A over number fields and describe the affine variant of Restriction of Scalars Chabauty. Let $K$ be a number field. Let $X$ be a smooth projective curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. Let $n=\#D(\overline{K})$. Let $S$ be a finite set of primes of $\mathcal{O}_{K}$ and let $\mathcal{X}$ be a regular model of $X$ over $\mathcal{O}_{K}$ that is $D$-transversal over primes in $S$. Fix a base point $P_{0}\in Y(K)$ and an $S$-integral reduction type $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ for $(\mathcal{X},\operatorname{\mathcal{D}})$ (see ˜3.8). We are interested in the set $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$ consisting of all $S$-integral points of reduction type $\Sigma$. Recall that $C(\Sigma)$ denotes the set of primes $\mathfrak{q}\in S$ for which $\Sigma_{\mathfrak{q}}$ is cuspidal, and $\Sigma^{{\mathrm{csp}}}\coloneqq(\Sigma_{\mathfrak{q}})_{\mathfrak{q}\in C(\Sigma)}$ is called the cuspidal part of $\Sigma$ (˜3.10).
5.1. Existence of Chabauty functions
Theorem 5.1.
With the notation as above, let $\mathfrak{p}\not\in S$ be a prime of $\mathcal{O}_{K}$. Assume
$r+\#C(\Sigma)+([K:\mathbb{Q}]-1)n<g+\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}+\#|D|+n_{2}(D)-1.$ (5.1)
Then there exist a non-zero log differential $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ and a constant $c\in K_{\mathfrak{p}}$, both depending on $\Sigma$, such that the function $\rho\colon\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}}$ given by
$\rho(P)\coloneqq\int_{P_{0}}^{P}\omega-c$ (5.2)
vanishes on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$. The differential $\omega$ can be chosen uniformly for all reduction types having the same cuspidal part.
Hypothesis (5.1) specialises to (1.2) for $K=\mathbb{Q}$, so ˜A is a special case of ˜5.1. Note that the function $\rho\colon\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}}$ in (5.2) has only finitely many zeros by §4.2.
Proof.
If $D$ is just a single $K$-rational point, Condition (5.1) becomes $r+\#C(\Sigma)<g$, and already the weaker condition $r<g$ is sufficient for the classical Chabauty method to apply, yielding a non-zero holomorphic differential $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1})$ such that $\int_{P_{0}}^{P}\omega=0$ for all rational points $P\in X(K)$. Hence, from now on assume $n>1$. In Section˜3.3 we constructed the Selmer set $\operatorname{Sel}(P_{0},\Sigma)\subseteq J_{Y}(K)$ with the property that the Abel–Jacobi map $\operatorname{AJ}_{P_{0}}\colon\mathcal{Y}(\mathcal{O}_{K,S})\to J_{Y}(K)$ maps all points of reduction type $\Sigma$ into $\operatorname{Sel}(P_{0},\Sigma)$ (see ˜3.13). The set $\operatorname{Sel}(P_{0},\Sigma)$ is the preimage under $\sigma\colon J_{Y}(K)\to V_{D}$ of a translate of a certain subgroup $U=U(\Sigma^{{\mathrm{csp}}})\subseteq V_{D}$. Let $W\coloneqq\sigma^{-1}(U)\subseteq J_{Y}(K)$, so that $\operatorname{Sel}(P_{0},\Sigma)$ is either empty or a translate $s_{0}+W$ of $W$. Consider the following diagram:
$\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$$\mathcal{Y}(\mathcal{O}_{K,S})$$\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$$s_{0}+W$$\operatorname{Sel}(P_{0},\Sigma)$$J_{Y}(K)$$J_{Y}(K_{\mathfrak{p}})$$\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\int_{P_{0}}}$$\scriptstyle{\log_{J_{Y}}}$ (5.3)
When $\operatorname{\mathrm{rk}}_{\mathbb{Z}}W$ is strictly less than $\dim_{K_{\mathfrak{p}}}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$, then the image of $W$ in $\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$ under the logarithm map is contained in a proper $K_{\mathfrak{p}}$-linear subspace, so there exists a non-zero $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ with $\langle\log_{J_{Y}}(w),\omega\rangle=0$ for all $w\in W$. Then, since $\operatorname{Sel}(P_{0},\Sigma)$ is a translate of $W$ (or empty), the map $s\mapsto\langle\log_{J_{Y}}(s),\omega\rangle$ is constant on $\operatorname{Sel}(P_{0},\Sigma)$. Pulling this back under the diagonal map in the diagram (5.3) shows that the map $P\mapsto\int_{P_{0}}^{P}\omega$ is constant on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$. By ˜3.14, the condition that the rank of $W$ be strictly less than the dimension of $\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ reads
$n_{1}(D)+n_{2}(D)-\#|D|-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}+r+\#C(\Sigma)<g+n-1,$
which, using $n_{1}(D)+2n_{2}(D)=[K:\mathbb{Q}]n$, simplifies to (5.1). ∎
Of course, when the base point $P_{0}$ is itself an $S$-integral point of reduction type $\Sigma$, then the constant $c$ in (5.2) is necessarily zero.
Remark 5.2.
The proof above shows that the differential $\omega$, the constant $c$ and hence the function $\rho$ in (5.2) depend on $\Sigma$ only through $\operatorname{Sel}(P_{0},\Sigma)$. That is, if $\operatorname{Sel}(P_{0},\Sigma)=\operatorname{Sel}(P_{0},\Sigma^{\prime})$ for two reduction types $\Sigma$ and $\Sigma^{\prime}$, then we get a single Chabauty function $\rho$ that vanishes on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}\cup\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma^{\prime}}$. The equality $\operatorname{Sel}(P_{0},\Sigma)=\operatorname{Sel}(P_{0},\Sigma^{\prime})$ is guaranteed if $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma)=\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma^{\prime})$ for all $\mathfrak{q}$.
5.2. Affine Restriction of Scalars Chabauty
Instead of locating $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$ for some suitable prime $\mathfrak{p}$ of $\mathcal{O}_{K}$ as in Section˜5.1, it can be advantageous to consider all primes $\mathfrak{p}$ of $\mathcal{O}_{K}$ lying over a fixed rational prime $p$ at once and try to locate $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$. This method is known as Restriction of Scalars Chabauty. It appeared first in [Sik13] in the context of rational points on projective curves.
The relevant diagram for Restrictions of Scalars Chabauty in our setting looks as follows:
$\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$$\mathcal{Y}(\mathcal{O}_{K,S})$$\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$$\operatorname{Sel}(P_{0},\Sigma)$$J_{Y}(K)$$\prod_{\mathfrak{p}|p}J_{Y}(K_{\mathfrak{p}})$$\prod_{\mathfrak{p}|p}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\prod_{\mathfrak{p}|p}\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\prod_{\mathfrak{p}|p}\int_{P_{0}}}$$\scriptstyle{\log_{J_{Y}}}$ (5.4)
Here, $p$ is a rational prime not divisible by any prime in $S$. Let $d\coloneqq[K:\mathbb{Q}]$. Since $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$ is now $d$-dimensional as a rigid analytic space over $\mathbb{Q}_{p}$, at least $d$ independent functions are needed to cut out a finite subset. The following theorem shows under which condition we find the right number of functions.
Theorem 5.3.
Let $p$ be a rational prime as above. Assume that the inequality
$r+\#C(\Sigma)\leq d(g-2)+n_{2}(D)+\#|D|+\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}$ (5.5)
is satisfied. Then there exist $d$ linearly independent (over $\mathbb{Q}_{p}$) tuples of logarithmic differentials $(\omega_{\mathfrak{p}}^{(k)})_{\mathfrak{p}}\in\prod_{\mathfrak{p}|p}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ and $p$-adic constants $c_{k}\in\mathbb{Q}_{p}$ ($k=1,\ldots,d$) such that the functions $\rho_{k}\colon\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to\mathbb{Q}_{p}$ given by
$\rho_{k}((P_{\mathfrak{p}})_{\mathfrak{p}|p})=\sum_{\mathfrak{p}|p}\mathrm{Tr}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}\left(\int_{P_{0}}^{P_{\mathfrak{p}}}\omega^{(k)}_{\mathfrak{p}}\right)-c_{k}$ (5.6)
vanish on the image of the diagonal embedding of $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$ in $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$.
Proof.
If $D$ consists just of a single $K$-rational point, Condition (5.5) is equivalent to $r+\#C(\Sigma)\leq d(g-1)$, and already the weaker condition $r\leq d(g-1)$ is enough to employ the classical Restriction of Scalars Chabauty method for the $K$-rational points on $X$. So we can assume that $n>1$. Consider the diagram (5.4). The Selmer set $\operatorname{Sel}(P_{0},\Sigma)$ constructed in Section˜3.3 is the preimage under $\sigma\colon J_{Y}(K)\to V_{D}$ of a translate of a certain subgroup $U=U(\Sigma^{{\mathrm{csp}}})\subseteq V_{D}$. Let $W\coloneqq\sigma^{-1}(U)\subseteq J_{Y}(K)$, so that $\operatorname{Sel}(P_{0},\Sigma)$ is either empty or a translate of $W$. Let $d^{\prime}$ be the codimension of the image of $W\otimes_{\mathbb{Z}}\mathbb{Q}_{p}$ in the $\mathbb{Q}_{p}$-vector space $\prod_{\mathfrak{p}|p}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$ along the bottom row of Diagram (5.4). Then there exist $d^{\prime}$ independent $\mathbb{Q}_{p}$-linear maps
$f_{k}\colon\prod_{\mathfrak{p}|p}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}\to\mathbb{Q}_{p}\qquad(k=1,\ldots,d^{\prime})$ (5.7)
vanishing on the image of $W$. For a finite-dimensional $K_{\mathfrak{p}}$-vector space $H$ one has a canonical isomorphism of $\mathbb{Q}_{p}$-vector spaces
$\operatorname{Res}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}(H^{\vee})\cong\operatorname{Res}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}(H)^{\vee}$
via $\phi\mapsto\mathrm{Tr}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}\circ\phi$. Here, $\operatorname{Res}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}$ means we view a $K_{\mathfrak{p}}$-vector space as a $\mathbb{Q}_{p}$-vector space via restriction of scalars, and $\mathrm{Tr}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}$ denotes the trace map $K_{\mathfrak{p}}\to\mathbb{Q}_{p}$. Applying this with $H=\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ and taking duals, one sees that each of the linear maps $f_{k}$ in (5.7) is given by
$f_{k}\colon(\phi_{\mathfrak{p}})_{\mathfrak{p}}\mapsto\sum_{\mathfrak{p}|p}(\mathrm{Tr}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}\circ\phi_{\mathfrak{p}})(\omega_{\mathfrak{p}}^{(k)})$
for a unique tuple of log differentials $(\omega_{\mathfrak{p}}^{(k)})_{\mathfrak{p}}\in\prod_{\mathfrak{p}|p}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$. Since $\operatorname{Sel}(P_{0},\Sigma)$ is a translate of $W$ (or empty), each map $f_{k}$ is constant on the Selmer set, say with value $c_{k}\in\mathbb{Q}_{p}$. Pulling $f_{k}$ back along the diagonal arrow in Diagram (5.4) results in the map
$\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to\mathbb{Q}_{p},\qquad(P_{\mathfrak{p}})_{\mathfrak{p}|p}\mapsto\sum_{\mathfrak{p}|p}\mathrm{Tr}_{K_{\mathfrak{p}}/\mathbb{Q}_{p}}\left(\int_{P_{0}}^{P_{\mathfrak{p}}}\omega^{(k)}_{\mathfrak{p}}\right)$
whose restriction to $\mathcal{Y}(\mathcal{O}_{K,S})$ is constant with value $c_{k}$ by commutativity of the diagram. Hence, we find $d^{\prime}$ many functions $\rho_{k}$ of the claimed form. It remains to see that $d^{\prime}\geq d$. By definition, $d^{\prime}$ can be bounded below by
$d(g+n-1)-\operatorname{\mathrm{rk}}W=d(g-1)-r+n_{2}(D)+\#|D|+\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}-\#C(\Sigma),$
where we used $n_{1}(D)+2n_{2}(D)=dn$ and ˜3.14. Now (5.5) ensures that $d^{\prime}\geq d$. ∎
Remark 5.4.
When applying ˜5.3 it may be convenient to choose a prime $p$ that is totally split in $K$, so that $K_{\mathfrak{p}}=\mathbb{Q}_{p}$ for all $\mathfrak{p}|p$. Then the integrals occurring in the Chabauty functions (5.6) are all defined over $\mathbb{Q}_{p}$ and the trace map can be omitted. Most implementations of Coleman integration in computer algebra systems are currently restricted to $\mathbb{Q}_{p}$.
Remark 5.5.
In the case that $\mathcal{Y}\colon y^{2}=f(x)$ with $f\in\mathcal{O}_{K}[x]$ monic, squarefree and of degree $2g+2$, condition (5.5) with $S=\emptyset$ specialises to $\operatorname{\mathrm{rk}}J(K)\leq dg-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}$. In this setting, functions $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to\mathbb{Q}_{p}$ vanishing on $\mathcal{Y}(\mathcal{O}_{K})$ are constructed in terms of $p$-adic heights in [GM25, Theorem 5.1].
Remark 5.6.
Since $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$ is a $d$-dimensional rigid analytic space, at least $d$ linearly independent Chabauty functions are needed to cut out a finite subset. However, having $d$ functions does not guarantee finiteness since the functions could conspire to have many zeros in common. The phenomenon of such unlikely intersections was first observed by Siksek [Sik13, §2] and studied in detail by Dogra [Dog24]. For affine curves of genus zero, limitations to the Restrictions of Scalars Chabauty–Skolem method (see ˜1.1) were studied in [Tri21].
6. Bounds on the number of $S$-integral points
Suppose we are in the setting of Section˜5 with $K=\mathbb{Q}$, so we have a curve $Y=X\smallsetminus D$ over $\mathbb{Q}$, a finite set of primes $S$, a regular model $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$ over $\mathbb{Z}$ that is $D$-transversal over primes in $S$, an auxiliary prime $p\not\in S$ and a base point $P_{0}\in Y(\mathbb{Q})$. If (1.1) holds, then ˜A applies for every $S$-integral reduction type $\Sigma$. Consequently, the set of $S$-integral points $\mathcal{Y}(\mathbb{Z}_{S})$ is contained in a finite union of zero loci of functions $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$ of the form (1.3). We give a bound on the number of zeros of such a function, extending the results of Coleman as outlined in [MP12, §5] for integrals of holomorphic differentials to those of logarithmic differentials. Following [MP12, Appendix A], we allow $p$ to be a prime of bad reduction.
6.1. Reduction of log differentials
We start by discussing the reduction of log differentials modulo $p$. This requires some care when $p$ is a prime of bad reduction since the sheaf of Kähler differentials on $\mathcal{X}_{\mathbb{Z}_{p}}$ fails to be a line bundle in this case and one has to pass through the canonical sheaf instead. We work in a general local setting.
As in §2.2, let $(R,\mathfrak{m},k)$ be a discrete valuation ring with uniformizer $\pi$ and fraction field $K$. Let $X/K$ be a smooth projective curve and let $\mathcal{X}/R$ be a regular model of $X$ with special fibre $\mathcal{X}_{\mathfrak{m}}$. Let $\emptyset\neq D\subseteq X$ be a finite set of closed points, $n=\#D(\overline{K})$, and $\operatorname{\mathcal{D}}$ the Zariski closure of $D$ in $\mathcal{X}$. Let $Y=X\smallsetminus D$ and $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. Let $0\neq\omega\in\mathrm{H}^{0}(X,\Omega^{1}(D))$ be a nonzero log differential. The canonical sheaf $\omega_{\mathcal{X}/R}$ is a line bundle on the model $\mathcal{X}$ whose restriction to the generic fibre agrees with the sheaf $\Omega^{1}_{X/K}$ of Kähler differentials, so we can view $\omega$ as a meromorphic section of $\omega_{\mathcal{X}/R}$ and let $\mathcal{K}=\mathrm{div}_{\mathcal{X}}(\omega)$ be the associated canonical divisor. Write $\mathcal{K}=\mathcal{H}+\mathcal{V}$ for a horizontal divisor $\mathcal{H}$ and a vertical divisor $\mathcal{V}$. Note that, if $\mathrm{div}(\omega)=\sum_{P\in|X|}a_{P}P$, then $\mathcal{H}=\sum_{P\in|X|}a_{P}\mathcal{P}$. In particular, since $\omega$ is a logarithmic differential, we have $a_{P}\geq 0$ for $P\in|Y|$ and $a_{P}\geq-1$ for $P\in D$, so $\mathcal{H}=\mathcal{H}_{e}-\mathcal{H}_{p}$ with $\mathcal{H}_{e}$ effective and $\mathcal{H}_{p}=\sum_{P\in D^{\prime}}\mathcal{P}$ with $D^{\prime}\subseteq D$ consisting of the poles of $\omega$.
Let $C$ be a component of $\mathcal{X}_{\mathfrak{m}}$ that occurs with multiplicity $1$ in $\mathrm{div}_{\mathcal{X}}(\pi)=\mathcal{X}_{\mathfrak{m}}$. Then we can multiply $\omega$ with an appropriate power of $\pi$ to obtain a log differential $\omega^{C}$ such that $C$ does not occur in $\mathcal{K}^{C}\coloneqq\mathrm{div}_{\mathcal{X}}(\omega^{C})$. We again write $\mathcal{K}^{C}=\mathcal{H}+\mathcal{V}^{C}$ with $\mathcal{H}$ the same horizontal divisor as before and $\mathcal{V}^{C}$ vertical. We reduce $\omega^{C}$ to $C^{\mathrm{sm}}=C\cap\mathcal{X}^{\mathrm{sm}}$ as follows: By removing all components and singular points of $\mathcal{X}_{\mathfrak{m}}$ except $C^{\mathrm{sm}}$, we obtain an open subscheme $\mathcal{X}^{C}\subseteq\mathcal{X}$. We may view $\omega^{C}$ as a (meromorphic) differential form on $\mathcal{X}^{C}$. As $C$ does not occur in $\mathcal{K}^{C}$, the restriction $\widetilde{\omega}^{C}\coloneqq\omega^{C}|_{C^{\mathrm{sm}}}$ is a nonzero (meromorphic) differential form on $C^{\mathrm{sm}}$. Note that the restriction of $\widetilde{\omega}^{C}$ to $C^{\mathrm{sm}}\cap\mathcal{Y}$ is holomorphic.
In conclusion, for every nonzero log differential $0\neq\omega\in\mathrm{H}^{0}(X,\Omega^{1}(D))$ and for every multiplicity-1 component $C$ of $\mathcal{X}_{\mathfrak{m}}$, there is a scalar multiple $\omega^{C}$ of $\omega$ which reduces to a nonzero meromorphic differential $\widetilde{\omega}^{C}$ on $C^{\mathrm{sm}}$ whose restriction to $C^{\mathrm{sm}}\cap\mathcal{Y}$ is holomorphic.
6.2. Bounding zeros of a log differential modulo $p$
Keep the notation from the previous subsection. We need two lemmas on the number of zeros of the reduction of a log differential. Given a multiplicity-1 component $C$ of $\mathcal{X}_{\mathfrak{m}}$ and given $\omega$ with reduction $\widetilde{\omega}^{C}$ on $C^{\mathrm{sm}}$, let
$n_{C}\coloneqq\sum_{\widetilde{R}\in(C^{\mathrm{sm}}\cap\mathcal{Y})(k)}\operatorname{ord}_{\widetilde{R}}(\widetilde{\omega}^{C}),$
so $n_{C}$ is the number of zeros of $\widetilde{\omega}^{C}$ in $(C^{\mathrm{sm}}\cap\mathcal{Y})(k)$ counted with multiplicities.
Lemma 6.1.
We have $n_{C}\leq\mathcal{H}_{e}\cdot C$.
Proof.
For $P\in(C^{\mathrm{sm}}\cap\mathcal{Y})(k)$ we have
$i_{P}(\mathcal{K}^{C},C)=i_{P}(\mathcal{H},C)=i_{P}(\mathcal{H}_{e},C)-i_{P}(\mathcal{H}_{p},C)=i_{P}(\mathcal{H}_{e},C),$
where the first equality is true because $P$ does not lie on any component of $\mathcal{V}^{C}$ as $C$ does not occur in $\mathcal{V}^{C}$, and the last equality is true because $P\in\mathcal{Y}(k)$ does not lie in the horizontal closure of a pole. As $\mathcal{H}_{e}$ is effective, summing over all $P$ yields
$n_{C}=\sum_{P\in(C^{\mathrm{sm}}\cap\mathcal{Y})(k)}i_{P}(\mathcal{K}^{C},C)\leq\sum_{P\in C}i_{P}(\mathcal{H}_{e},C)[k(P):k]=\mathcal{H}_{e}\cdot C.\qed$
Lemma 6.2.
With the above notation, we have
$\sum_{C\textnormal{ of multiplicity }1}n_{C}\leq 2g-2+n.$
Proof.
The divisor $\mathcal{H}_{e}$ is effective, hence $\mathcal{H}_{e}\cdot C\geq 0$ for every component $C$ of $\mathcal{X}_{\mathfrak{m}}=\sum_{C}m_{C}C$. By Lemma˜6.1, we have
[…]
Here $\mathcal{K}\cdot\mathcal{X}_{\mathfrak{m}}=2g-2$ by the adjunction formula [Liu06, Proposition 9.1.35], and $\mathcal{V}\cdot\mathcal{X}_{\mathfrak{m}}=0$ as $\mathcal{V}$ is vertical [Liu06, Proposition 9.1.21(a)]. Finally,
$\mathcal{H}_{p}\cdot\mathcal{X}_{\mathfrak{m}}=\sum_{P\in D^{\prime}}\mathcal{P}\cdot\mathcal{X}_{\mathfrak{m}}=\sum_{P\in D^{\prime}}[k(P):K]\leq\sum_{P\in D}[k(P):K]=n$
by [Liu06, Proposition 9.1.30], which finishes the proof. ∎
6.3. Bounding the number of zeros of a logarithmic abelian integral
We stay in the local setting as above but specialise to the case $R=\mathbb{Z}_{p}$, so $Y=X\smallsetminus D$ is a curve over $\mathbb{Q}_{p}$ with regular model $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$ over $\mathbb{Z}_{p}$, and $\omega\in\mathrm{H}^{0}(X,\Omega^{1}(D))$ is a nonzero log differential. Fix a base point $P_{0}\in Y(\mathbb{Q}_{p})$. We prove a bound on the number of points $P$ in a residue disc for which $\int_{P_{0}}^{P}\omega$ is equal to a prescribed constant. We need the following lemma, which is taken from [MP12, Lemma 5.1] and proved by looking at the Newton polygon.
Lemma 6.3.
[…]
Proposition 6.4.
Let $C\subseteq\mathcal{X}_{p}^{\mathrm{sm}}$ be a multiplicity-1 component, let $\widetilde{R}\in(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})$ and let $m=\operatorname{ord}_{\widetilde{R}}(\widetilde{\omega}^{C})$. Assume m<p2m<p-2. Then, for any $c\in\mathbb{Q}_{p}$, the number of $P\in\mathcal{Y}(\mathbb{Z}_{p})$ reducing to $\widetilde{R}$ and satisfying $\int_{P_{0}}^{P}\omega=c$ is at most $m+1$.
Proof.
We may rescale $\omega$ (and $c$) and assume $\omega=\omega^{C}$. Note that $m\geq 0$ as $\widetilde{R}$ is not contained in the support of $\operatorname{\mathcal{D}}_{p}$. By Hensel’s lemma, $\widetilde{R}$ lifts to a point $R\in\mathcal{Y}(\mathbb{Z}_{p})$. Let $t$ be a rational function on $X$ that reduces to a uniformiser on $\mathcal{X}_{p}$ at $\widetilde{R}$. The function $t$ maps the residue disc of $\widetilde{R}$ bijectively to $p\mathbb{Z}_{p}$. The order of vanishing of $f^{\prime}(t)~(\operatorname{mod}p)$ at $t=0$ is equal to $m$ and the result follows from Lemma˜6.3. ∎
By summing over all residue discs of $\operatorname{\mathbb{F}}_{p}$-points lying on a fixed component we obtain the following.
Proposition 6.5.
Let $C\subseteq\mathcal{X}_{p}^{\mathrm{sm}}$ be a multiplicity-1 component and let $n_{C}$ be the number of zeros of $\widetilde{\omega}^{C}$ in $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})$ counted with multiplicities. Assume $n_{C}<p-2$. Then, for any $c\in\mathbb{Q}_{p}$, the number of $P\in\mathcal{Y}(\mathbb{Z}_{p})$ reducing onto $C$ and satisfying $\int_{P_{0}}^{P}\omega=c$ is at most $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})+n_{C}$.
Proof.
Every point $P\in\mathcal{Y}(\mathbb{Z}_{p})$ reducing onto $C$ reduces to a unique $\widetilde{R}\in(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})$. Let $m_{\widetilde{R}}=\operatorname{ord}_{\widetilde{R}}(\widetilde{\omega}^{C})$, so that $n_{C}=\sum m_{\widetilde{R}}$. The assumption on $p$ implies $m_{\widetilde{R}}<p-2$, so ˜6.4 implies that the number of $P\in\mathcal{Y}(\mathbb{Z}_{p})$ reducing to $C$ and satisfying $\int_{P_{0}}^{P}\omega=c$ is equal to
$\sum_{\widetilde{R}\in(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})}\#\left\{P\in\mathcal{Y}(\mathbb{Z}_{p})\;\Big|\;{\mathrm{red}}_{p}(P)=\widetilde{R}\text{ and }\int_{P_{0}}^{P}\omega=c\right\}\leq\sum_{\widetilde{R}\in(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})}(m_{\widetilde{R}}+1),$
which is equal to $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})+n_{C}$. ∎
6.4. Effective bounds on the number of $S$-integral points
We return to the global situation, so $Y=X\smallsetminus D$ is an affine curve over a number field $K$ with regular model $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$ over $\mathcal{O}_{K}$ which is $D$-transversal over primes in $S$. We use the results of the previous subsections to prove a Coleman bound on the number of $S$-integral points.
Theorem 6.6.
Let $\Sigma$ be an $S$-integral reduction type such that the Affine Chabauty Condition (5.1) is satisfied. Let $\mathfrak{p}\not\in S$ be a prime with $K_{\mathfrak{p}}=\mathbb{Q}_{p}$ and assume $p>2g+n$. Let $C\subseteq\mathcal{X}_{\mathfrak{p}}$ be the component of the mod-$\mathfrak{p}$ fibre determined by the reduction type $\Sigma$. Then
$\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}\leq\#(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})+2g-2+n.$
Proof.
By ˜5.1, there exist a non-zero log differential $\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ and a constant $c$ such that $\int_{P_{0}}^{P}\omega=c$ for all $P\in\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$. Hence, we have the inclusion
$\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}\subseteq\left\{P\in\mathcal{Y}(\mathbb{Z}_{p})\;\Big|\;{\mathrm{red}}_{p}(P)\in C\text{ and }\int_{P_{0}}^{P}\omega=c\right\}.$
We will bound the size of the latter set using ˜6.5. Rescaling $\omega$ and $c$ by a power of $p$ yields a log differential $\omega^{C}$ and a constant $c^{C}$ such that the restriction $\widetilde{\omega}^{C}\coloneqq\omega^{C}|_{C^{\mathrm{sm}}}$ is defined and nonzero. Let $n_{C}$ be the number of zeros of $\widetilde{\omega}^{C}$ in $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})$ counted with multiplicity. Lemma˜6.2 implies $n_{C}\leq 2g-2+n$, and this is <p2<p-2 by the assumption on $p$. So ˜6.5 applies and yields
$\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}\leq\#(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})+n_{C}.$ (6.1)
The claim now follows from $n_{C}\leq 2g-2+n$. ∎
Proof of ˜C.
The idea is to bound $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ via (6.1) for each $S$-integral reduction type and sum over all $\Sigma$. Let us first consider the situation with fixed prime-to-$p$ reduction type $\Sigma^{\prime}\in\prod_{\ell\neq p}\mathcal{C}_{\ell}$. Let $\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma^{\prime}}$ be the set of $S$-integral points whose prime-to-$p$ reduction type is equal to $\Sigma^{\prime}$. For a multiplicity-1 component $C\subseteq\mathcal{X}_{p}$ with $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})\neq\emptyset$, write $\Sigma=(\Sigma^{\prime},C)$ for the reduction type whose prime-to-$p$ part is $\Sigma^{\prime}$ and whose $p$-part is $\Sigma_{p}=C^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}$. Since $p\not\in S$, the cuspidal part of $\Sigma$ (˜3.10) is determined by $\Sigma^{\prime}$. The Affine Chabauty Condition (1.1) implies (5.1), so by ˜5.1 there exist a non-zero log differential $\omega=\omega(\Sigma^{\prime})\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))$ and constants $c_{\Sigma}\in\mathbb{Q}_{p}$ such that $\int_{P_{0}}^{P}\omega=c_{\Sigma}$ for all $P\in\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}$ of reduction type $\Sigma=(\Sigma^{\prime},C)$. Let $n_{C}$ be the number of zeros of $\widetilde{\omega}^{C}$ in $(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})$ counted with multiplicity, so we have
$\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma}\leq\#(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})+n_{C}$
for all $\Sigma=(\Sigma^{\prime},C)$ by (6.1). Summing over all $C$ and using Lemma˜6.2 we obtain
$\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma^{\prime}}\leq\sum_{C}(\#(C^{\mathrm{sm}}\cap\mathcal{Y})(\operatorname{\mathbb{F}}_{p})+n_{C})\leq\#\mathcal{Y}^{\mathrm{sm}}_{p}(\operatorname{\mathbb{F}}_{p})+2g-2+n.$ (6.2)
Finally, since $\mathcal{Y}(\mathbb{Z}_{S})=\sum_{\Sigma^{\prime}}\#\mathcal{Y}(\mathbb{Z}_{S})_{\Sigma^{\prime}}$, a bound on $\mathcal{Y}(\mathbb{Z}_{S})$ is obtained by multiplying the right hand side of (6.2) by the number of all possible prime-to-$p$ reduction parts, which is
$\prod_{\ell\neq p}\#\mathcal{C}_{\ell}=\prod_{\ell\in S}(\#\mathcal{C}_{\ell}^{0}+\#\mathcal{C}_{\ell}^{1})\prod_{\ell\not\in S\cup\{p\}}\#\mathcal{C}_{\ell}^{0}=\prod_{\ell\in S}(n_{\ell}+\#(\mathcal{X}_{{\ell}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell}))\prod_{\ell\not\in S\cup\{p\}}n_{\ell}.\qed$
Remark 6.7.
The bound in ˜C can be improved by observing that different reduction types can lead to the same combination $(\omega,c)$ and hence to the same Chabauty function $\rho$ as in (1.3). Namely, let $T$ denote the set of all primes $\ell$ for which the horizontal divisor $\operatorname{\mathcal{D}}$ intersects $\mathcal{X}_{\ell}$ in more than one component. Note that $T$ is a subset of the set of primes of bad reduction for $\mathcal{X}$. In particular, $T$ is finite. We can assume that $(\mathcal{X}_{\ell}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell})\neq\emptyset$ for all $\ell\in S$. If two $S$-integral reduction types $\Sigma$ and $\Sigma^{\prime}$ have the same cuspidal part and satisfy $\Sigma_{\ell}=\Sigma^{\prime}_{\ell}$ for every $\ell\in T$, then they give rise to the same pair $(\omega,c)$. Indeed, by ˜2.16 we then have $\mathfrak{S}_{\ell}(P_{0},\Sigma_{\ell})=\mathfrak{S}_{\ell}(P_{0},\Sigma^{\prime}_{\ell})$ for all primes $\ell$, so we get the same $(\omega,c)$ by ˜5.2. Moreover, by ˜2.31, at primes in $S\smallsetminus T$ any cuspidal reduction type automatically covers all component reduction types, so it is enough to consider $\Sigma$ with $S\smallsetminus T\subseteq C(\Sigma)$. Moreover, by the same remark, at primes $\ell$ in $S\cap T$ only those components not containing a point in $(\mathcal{X}_{\ell}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell})$ need to be considered separately. This shows that the bound of ˜C can be improved to
$\#\mathcal{Y}(\mathbb{Z}_{S})\leq(\#\mathcal{Y}^{\mathrm{sm}}_{p}(\operatorname{\mathbb{F}}_{p})+2g-2+n)\sum_{S\smallsetminus T\subseteq S_{0}\subseteq S}\;\prod_{\ell\in S_{0}}\#(\mathcal{X}_{\ell}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\ell})\prod_{\ell\in T\smallsetminus(S_{0}\cup\{p\})}n_{\ell}^{\prime}$ (6.3)
where we set $n_{\ell}^{\prime}=n_{\ell}$ for $\ell\not\in S$ but define $n_{\ell}^{\prime}$ as the number of components of $\mathcal{X}_{\ell}$ containing at least one smooth $\operatorname{\mathbb{F}}_{\ell}$-point but no smooth $\operatorname{\mathbb{F}}_{\ell}$-point on $\operatorname{\mathcal{D}}$ for $\ell\in S$. In the case $S=\emptyset$ this simplifies to
$\mathcal{Y}(\mathbb{Z})\leq(\#\mathcal{Y}^{\mathrm{sm}}_{p}(\operatorname{\mathbb{F}}_{p})+2g-2+n)\prod_{\ell\in T\smallsetminus\{p\}}n_{\ell}^{\prime}.$ (6.4)
We will use the refined bound (6.4) in ˜7.1.
Example 6.8.
Consider the affine curve $\mathcal{Y}\subseteq\mathbb{A}^{2}_{\mathbb{Z}}$ defined by the equation
$y^{3}=x^{3}+7x^{2}+x.$
Our results can be used to bound the number of solutions to this equation in $\mathbb{Z}[1/q]$ for any fixed prime $q$. A smooth compactification of $Y=\mathcal{Y}_{\mathbb{Q}}$ is the rank 1 elliptic curve $X:Y^{\prime 2}+7Y^{\prime}Z^{\prime 3}=X^{\prime 3}-Z^{\prime 3}$, into which $Y$ embeds via $(x,y)\mapsto[y:1:x]$. Let $\mathcal{X}/\mathbb{Z}$ be the Weierstrass model defined by the same equation. This model is regular, the primes of bad reduction are $3$ and $5$, and both bad fibres have only one component. In the coordinates $x^{\prime}=X^{\prime}/Z^{\prime}$ and $y^{\prime}=Y^{\prime}/Z^{\prime}$, the divisor of cusps in $X$ is $D=\{y^{\prime}=0\}$, consisting of the rational point $Q_{1}=(1,0)$ and the quadratic point $Q_{2}$ with $Q_{2}({\overline{\mathbb{Q}}})=\{(\zeta_{3},0),(\zeta_{3}^{-1},0)\}$, where $\zeta_{3}$ is a primitive third root of unity. The horizontal closures $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are normal and only intersect in the fibre over $3$, so the model $\mathcal{X}$ is $D$-transversal at all primes except $3$. We take $S=\{q\}$ for a prime $q\neq 3$. (In order to treat the case $q=3$ one could blow up at the point in the mod-3 fibre where $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ meet and work with the resulting $D$-transversal model.) Let $\Sigma$ be an $S$-integral reduction type. By ˜2.31 it is enough to consider reduction types with $\Sigma_{q}$ cuspidal. The number of such reduction types is $\operatorname{\mathcal{D}}(\operatorname{\mathbb{F}}_{q})$, which is 1 or 3, depending on whether $q$ splits in $\mathbb{Q}(\zeta_{3})$. The Affine Chabauty Condition (1.2) is satisfied:
$r+\#C(\Sigma)=1+1<1+1+1-1=g+\#|D|+n_{2}(D)-1.$
Thus, by ˜6.6, there are at most $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})+3$ $S$-integral points of reduction type $\Sigma$, for any prime $p>2g+n=5$ with $p\neq q$. Assume that $q\neq 7$, then taking $p=7$, we have $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})=3$, so we obtain the bound $\mathcal{Y}(\mathbb{Z}[1/q])_{\Sigma}\leq 6$. Multiplying with the number of reduction types which are cuspidal at $q$, we obtain
$\mathcal{Y}(\mathbb{Z}[\tfrac{1}{q}])\leq\begin{cases}6,&\text{if $q\equiv-1\bmod 3$},\\ 18,&\text{if $q\equiv+1\bmod 3$}.\end{cases}$
7. Application: even degree hyperelliptic curves
7.1. Recovering Linear Quadratic Chabauty
Let $\mathcal{Y}/\mathbb{Z}$ be a hyperelliptic curve defined by an affine equation $y^{2}=f(x)$ with $f\in\mathbb{Z}[x]$ squarefree of degree $2g+2$. Assume that the leading coefficient of $f$ is a square in $\mathbb{Z}$, so that $Y=\mathcal{Y}_{\mathbb{Q}}$ has two rational cusps at infinity. Suppose that the Mordell–Weil rank $r$ equals $g$. Let $p$ be a prime of good reduction and fix $P_{0}\in\mathcal{Y}(\mathbb{Z})$ for simplicity. In [GM25], the authors use $p$-adic heights to construct a nontrivial function $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$ and a finite set $T$ such that $\rho(\mathcal{Y}(\mathbb{Z}))\subseteq T$. On the other hand, our ˜5.1 (applied to a desingularised model, see §3.4) shows the existence of a log differential $0\neq\omega\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1}(D))$ and finitely many constants $c_{\Sigma}$ indexed by integral reduction types such that $\int_{P_{0}}^{P}\omega=c_{\Sigma}$ holds for all $P\in\mathcal{Y}(\mathbb{Z})_{\Sigma}$. It is natural to ask whether we get the same functions, i.e. whether the function $\rho$ of [GM25] is given by the integral of a log differential. This is indeed the case and becomes clear once we recall the construction of $\rho$.
For each prime $\ell$ we have a local Coleman–Gross height pairing $h_{\ell}$ defined on degree zero divisors on $X_{\mathbb{Q}_{\ell}}$ with disjoint support, taking values in $\mathbb{Q}_{p}$ [CG89]. For $\ell=p$ the pairing depends on a choice of subspace $W\subseteq\mathrm{H}^{1}_{\operatorname{dR}}(X/\mathbb{Q}_{p})$ which is complementary to the space of holomorphic differentials and isotropic with respect to the cup product pairing. The choice is used to associate to each degree zero divisor $E$ on $X_{\mathbb{Q}_{p}}$ a uniquely determined meromorphic differential $\omega_{E}$ with at most simple poles and with residue divisor equal to $E$. The local height pairing at $p$ is then defined as
$h_{p}(E,F)\coloneqq\int_{F}\omega_{E}.$
For $\ell\neq p$ the height pairing is defined in terms of intersection theory on a regular model of $X$ over $\mathbb{Z}_{\ell}$ by
$h_{\ell}(E,F)\coloneqq-\log(\ell)\cdot i_{\ell}(\mathcal{E},\Psi_{\ell}(F))$
where $\mathcal{E}$ is the horizontal extension of $E$ and $\Psi_{\ell}(F)=\mathcal{F}+\Phi_{\ell}(F)$ is as in §2.3. The global Coleman–Gross height pairing is then defined for degree zero divisors with disjoint support on $X$ as the sum of the local height pairings over all places:
$h(E,F)\coloneqq\sum_{\ell}h_{\ell}(E,F).$
The global height pairing vanishes when one of the divisors is principal, so it is well-defined on the Jacobian of $X$. This is not true for the local height pairings, although one can show that when $E$ is supported on cusps, the function $F\mapsto h_{\ell}(E,F)$ is well-defined on the generalised Jacobian $J_{Y}$.
On a hyperelliptic curve $y^{2}=f(x)$ with $\deg(f)=2g+2$ as above we have two $\mathbb{Q}$-rational points at infinity. Denote them by $\infty_{+}$ and $\infty_{-}$ and fix the degree zero divisor $E\coloneqq\infty_{+}$-$\infty_{-}$ as first argument of the height pairings. The function $h(E,-)\colon J(\mathbb{Q})\to\mathbb{Q}_{p}$ is a well-defined homomorphism on the Jacobian. One may assume that the closure of $J(\mathbb{Q})$ has finite index in $J(\mathbb{Q}_{p})$, for otherwise one can find a non-zero holomorphic differential whose integral vanishes on all of $X(\mathbb{Q})$ by the classical Chabauty method. When the closure of $J(\mathbb{Q})$ has finite index in $J(\mathbb{Q}_{p})$, every homomorphism $J(\mathbb{Q})\to\mathbb{Q}_{p}$ is given by the integral of a holomorphic differential, in particular there exists $\omega_{\operatorname{hol}}\in\mathrm{H}^{0}(X_{\mathbb{Q}_{p}},\Omega^{1})$ such that
[…]
At $\ell=p$, the local height $h_{p}(\infty_{+}-\infty_{-},-)$ is given by the integral of a logarithmic differential $\omega_{E}$ with residues $1$ and 1-1 at $\infty_{+}$ and $\infty_{-}$, respectively. Subtracting the local height at $p$ from the global height and taking $F\coloneqq P-P_{0}$, we have
$\int_{P_{0}}^{P}\omega_{\operatorname{hol}}-\int_{P_{0}}^{P}\omega_{E}=\sum_{\ell\neq p}h_{\ell}(\infty_{+}-\infty_{-},P-P_{0})\quad\text{for $P\in Y(\mathbb{Q})$}.$
The expression on the left hand side defines a function $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$, whereas the right hand side takes only finitely many values for $P\in\mathcal{Y}(\mathbb{Z})$, corresponding to the possible reduction types of $P$.
The function $\rho$ obtained in this way is given as the integral of the log differential $\omega\coloneqq\omega_{\operatorname{hol}}-\omega_{E}$ and thus of the same form as the function obtained from our ˜5.1. Apart from working also for primes $p$ of bad reduction, our approach has the advantage that it yields directly the log differential $\omega$ without the need to determine $\omega_{\operatorname{hol}}$ and $\omega_{E}$ individually. Note that the splitting of $\omega$ into $\omega_{\operatorname{hol}}$ and $\omega_{E}$ depends on the choice of complementary subspace $W\subseteq\mathrm{H}^{1}_{\operatorname{dR}}(X/\mathbb{Q}_{p})$ but $\omega$ itself is unique up to scaling, independently of any choices.
The height-based approach can be generalised to more general affine curves admitting a divisor of degree zero supported in the cusps. This is satisfied whenever $D$ contains at least two closed points, and in this case there are #|D|1\#|D|-1 independent such divisors. On the other hand, our Affine Chabauty Condition (1.1) in the case $r=g$ and $S=\emptyset$ reads $0<\#|D|+n_{2}(D)-1$, so our approach yields more independent Chabauty functions in general and also applies when there is just a single cusp with non-totally real residue field, for example for even-degree hyperelliptic curves $y^{2}=f(x)$ where $f$ has negative leading coefficient. One reason for this is that the local height pairing $h_{\ell}$ at a prime $\ell\neq p$ only captures a total intersection number $i_{\ell}(-,-)$ in the mod-$\ell$ fibre of a regular model, whereas our $D$-intersection map captures the intersection numbers $i_{x}(-,-)$ at individual closed points $x$, which by (2.2) is more refined information. It seems likely that the height-based method corresponds to the Chabauty–Kim method for the quotient of the abelianised fundamental group of $Y$ described in [LLM23, Remark 4.3], which is an extension of the Tate module of $J$ by a Tate representation, whereas our method corresponds to the full abelianisation, which is an extension of the Tate module by an Artin–Tate representation. In line with Dogra’s terminology [Dog25], one might call the weaker variant “skimmed linear Chabauty”.
On the other hand, using the methods of this paper we cannot handle integral points on hyperelliptic curves of odd degree with $r=g$ as in [BBM16]. In this case there is only one point at infinity and the proof of Theorem 3.1 of loc. cit. uses the function $x\mapsto h(x-\infty,x-\infty)$ in which both arguments of the height pairing depend on $x$, leading to functions involving not just single but double Coleman integrals. An “Affine Quadratic Chabauty” method should be able to recover the results of [BBM16] as a special case.
7.2. A bound for integral points on hyperelliptic curves over $\mathbb{Q}$
Consider a hyperelliptic curve $\mathcal{Y}/\mathbb{Z}$ defined by an affine equation $y^{2}=f(x)$ with $f\in\mathbb{Z}[x]$ monic and squarefree of degree $2g+2$ and suppose that the Mordell–Weil rank $r$ equals $g$. As discussed in §7.1, we obtain a non-zero log differential whose integral takes only finitely many values on integral points. We show that in many cases it takes just one value. With the results from §6 this leads to a good bound on $\mathcal{Y}(\mathbb{Z})$.
Theorem 7.1.
Let $f=x^{2g+2}+f_{2g+1}x^{2g+1}+\dots+f_{0}\in\mathbb{Z}[x]$ be squarefree with $2g+2\geq 4$. Assume that the reduction of $f$ modulo every odd prime is not a square. Assume that either the coefficient $f_{2g+1}$ is odd, or that $f=4P+Q^{2}$ with $P,Q\in\mathbb{Z}[x]$ and $Q$ monic of degree $g+1$ such that the reduction of $y^{2}+Q(x)y-P(x)\in\mathbb{Z}[x][y]$ modulo $2$ does not have a zero $y_{0}\in\operatorname{\mathbb{F}}_{2}[x]$ with $\deg(y_{0})\leq g+1$. Let $p>2g+2$ be a prime such that the reduction of $f$ modulo $p$ remains separable. Assume that the hyperelliptic curve given by $y^{2}=f(x)$ over $\mathbb{Q}$ has Mordell–Weil rank equal to $g$. Then the affine curve
$\mathcal{Y}\colon\quad y^{2}=f(x)$ (7.1)
has at most $\mathcal{Y}(\operatorname{\mathbb{F}}_{p})+2g$ integral points.
We remark that these technical assumptions on $f$ modulo all primes can be checked very quickly.
Proof.
Let $\mathcal{X}/\mathbb{Z}$ be the projective scheme given by the affine equation (7.1) as in [Liu23, Definition 1.1], i.e. we work in weighted projective space with weights $(1,g+1,1)$. Let $X$ be its generic fibre and let $D=\{\infty_{+},\infty_{-}\}$ consist of the two points $\infty_{\pm}\coloneq(1:\pm 1:0)$ at infinity, which are $\mathbb{Q}$-rational. Let $Y=X\smallsetminus D$ and $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$ as usual. We are interested in $\mathcal{Y}(\mathbb{Z})$. Assume there exists a rational point $P_{0}\in Y(\mathbb{Q})$, otherwise there is nothing to prove. Then $Y$ satisfies (1.1) for empty $S$, but we cannot directly apply ˜C as $\mathcal{X}$ might not be regular. Instead, we need to work with a desingularisation of $\mathcal{X}$ as in §3.4. The assumptions on $f$ imply that
  1. (1)
    $\mathcal{X}_{p}$ is smooth;
  2. (2)
    for odd $\ell$, the special fibre $\mathcal{X}_{\ell}$ is irreducible and $\infty_{\pm}$ reduce to smooth points of $\mathcal{X}_{\ell}$, and the scheme $\mathcal{X}$ is normal at the points of $\mathcal{X}_{\ell}$ by [Liu23, Lemma 2.2]. In this case, we say $\mathcal{X}$ satisfies $(\ast)$ at $\ell$;
  3. (3)
    if $f_{2g+1}$ is odd, then $\mathcal{X}$ satisfies $(\ast)$ at $2$ by [Liu23, Lemma 2.3(1)(c)];
  4. (4)
    otherwise, let $\mathcal{X}^{\prime}$ be given by the affine equation
    $\mathcal{X}^{\prime}\colon\quad y^{2}+Q(x)y=P(x)$
    and $\mathcal{X}^{\prime}\to\mathcal{X}$ be given by $(x,y)\mapsto(x,2y+Q(x))$. Then $\mathcal{X}^{\prime}$ satisfies $(\ast)$ at $2$ by [Liu23, Lemma 2.3(1)(a)].
Let $\mathcal{X}^{\prime}$ be the normalisation of $\mathcal{X}$. Then $\mathcal{X}^{\prime}$ satisfies $(\ast)$ at all primes $\ell$. Now let $\mathcal{X}^{\prime\prime}$ be the strong desingularisation of $\mathcal{X}^{\prime}$ as in [Liu06, Theorem 9.3.44]. We have $\mathcal{Y}(\mathbb{Z})\subseteq\mathcal{Y}^{\prime\prime}(\mathbb{Z})$ and $\mathcal{Y}_{p}\cong\mathcal{Y}^{\prime\prime}_{p}$ by (1). As $\mathcal{X}^{\prime\prime}\to\mathcal{X}^{\prime}$ is an isomorphism away from the singular locus, we see that $\infty_{\pm}$ reduce to the same component of $\mathcal{X}^{\prime\prime}_{\ell}$ at every prime $\ell$. So $T$ is empty in the notation of ˜6.7 and the bound (6.4) applied to $\mathcal{Y}^{\prime\prime}$ yields the result. ∎
Remark 7.2.
The condition on $p$ in ˜7.1 can be slightly weakened: instead of requiring the WeierstraSS model $\mathcal{X}$ defined by $y^{2}=f(x)$ to have good reduction at $p$ it is enough that $f$ is not a square modulo $p$ and $\mathcal{X}$ is regular at all $\operatorname{\mathbb{F}}_{p}$-points in $\mathcal{Y}$. One then has the bound $\mathcal{Y}(\mathbb{Z})\leq\#\mathcal{Y}_{p}^{\mathrm{sm}}(\operatorname{\mathbb{F}}_{p})+2g$.
The bound of ˜7.1 is optimal, as the following examples demonstrate.
Example 7.3.
[…]
Example 7.4.
($g=2$) Let $f(x)=x^{6}-28x^{2}+4$. It is square-free, has discriminant $2^{16}\cdot 43^{2}\cdot 127^{2}$ and $f$ is not a square modulo $43$ or $127$. One readily checks the condition at $2$ with $Q=x^{3}$ and $P=1-7x^{2}$. The hyperelliptic curve defined by $y^{2}=f(x)$ is bielliptic, hence its Jacobian $J$ is isogenous to the product of the two elliptic curves $y^{2}=x^{3}-28x+4$ and $y^{2}=x^{3}-28x^{2}+16$ [BP24, p.216]. These are isomorphic to the elliptic curves with LMFDB labels 43688.a1 and 5461.a1 respectively, which both have rank $1$, so $J(\mathbb{Q})$ has rank $2$. Take $p=7$. Then $\mathcal{Y}\colon y^{2}=f(x)$ satisfies $\mathcal{Y}(\operatorname{\mathbb{F}}_{7})=2$, hence ˜7.1 gives the bound $\mathcal{Y}(\mathbb{Z})\leq 6$. A quick search finds the six points $(0,\pm 2)$, $(7,\pm 341)$, $(-7,\pm 341)$, thus $\mathcal{Y}(\mathbb{Z})=\{(0,\pm 2),(7,\pm 341),(-7,\pm 341)\}$.
7.3. Hyperelliptic curves over imaginary quadratic fields
Let $K$ be an imaginary quadratic field with ring of integers $\mathcal{O}_{K}$, and let $Y/K$ be an affine hyperelliptic curve defined by
$Y\colon\quad y^{2}=f(x)$ (7.2)
where $f(x)\in\mathcal{O}_{K}[x]$ is a squarefree polynomial of degree $2g+2\geq 4$ whose leading coefficient is a square in $\mathcal{O}_{K}$. Let $\mathcal{Y}\subseteq\mathbb{A}^{2}_{\mathcal{O}_{K}}$ be the model over $\mathcal{O}_{K}$ defined by the same equation. Write $Y=X\smallsetminus D$ and $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$ with $X$ and $\mathcal{X}$ given by taking the closure in the weighted projective space $\mathbb{P}(1,g+1,1)$. Assume that $r=\operatorname{\mathrm{rk}}J(K)$ is equal to  $g$. Since the degree of $f$ is even and the leading coefficient is a square, there are two points at infinity, which are both $K$-rational. Thus (5.1) is satisfied for empty $S$ and we may apply ˜5.1 to determine the $\mathcal{O}_{K}$-points of $\mathcal{Y}$. Note that the model $\mathcal{X}/\mathcal{O}_{K}$ defined by the equation (7.2) might not be regular, in which case we have to pass to a desingularisation $\mathcal{X}^{\prime}\to\mathcal{X}$ in order to specify the reduction types as explained in Section˜3.4. Here we only use a single prime $\mathfrak{p}$; one could also use all primes lying over a fixed rational prime as in ˜5.3 and [GM25, §5].
As a concrete illustration of our ˜5.1, we determine the $\mathbb{Z}[\zeta_{3}]$-points on the genus 2 hyperelliptic curve with LMFDB label 1549.a.1549.1 defined by
$Y\colon\quad y^{2}=x^{6}-4x^{5}+2x^{4}+6x^{3}+x^{2}-10x+1.$ (7.3)
Theorem 7.5.
The set $\mathcal{Y}(\mathbb{Z}[\zeta_{3}])$ is equal to
$\left\{(0,\pm 1),\;(2,\pm 1),\;(1,\pm\sqrt{-3})\right\}.$ (7.4)
It is convenient to choose an auxiliary prime $\mathfrak{p}$ with $K_{\mathfrak{p}}=\mathbb{Q}_{p}$. We choose $\mathfrak{p}\coloneqq(2-\sqrt{-3})$ which divides the rational prime $7$. The Jacobian has rank 1 over $\mathbb{Q}$ but rank $2$ over $K=\mathbb{Q}(\zeta_{3})$. For the lower bound $\operatorname{\mathrm{rk}}J(K)\geq 2$ it suffices to check that
$\det\begin{pmatrix}\int_{P_{0}}^{P_{1}}\omega_{0}&\int_{P_{0}}^{P_{1}}\omega_{1}\\[5.69054pt] \int_{P_{0}}^{P_{2}}\omega_{0}&\int_{P_{0}}^{P_{2}}\omega_{1}\end{pmatrix}\neq 0,$
where $P_{0}=(0,1)$, $P_{1}=(2,1)$, $P_{2}=(1,\sqrt{-3})$, and $\omega_{j}=x^{j}\frac{\mathrm{d}x}{y}$. The non-vanishing of the determinant implies that the divisors $P_{1}]-[P_{0}]$ and $P_{2}]-[P_{0}]$ are linearly independent in the Jacobian. The upper bound $\operatorname{\mathrm{rk}}J(K)\leq 2$ can be obtained by an application of the RankBound function in Magma. The model $\mathcal{X}_{0}$ over $\mathbb{Z}$ given by (7.3) has bad reduction at 2 and 1549. The model $\mathcal{X}^{\prime}_{0}$ with the minimal equation
$y^{2}+(x^{3}+x+1)y=-x^{5}+x^{3}-3x$
has good reduction at 2 and the map $\mathcal{X}_{0}^{\prime}\to\mathcal{X}_{0}$, $(x,y)\mapsto(x,2y+x^{3}+x+1)$ is an isomorphism over all odd primes, so the only bad prime for this model is $q=1549$. An application of the RegularModel function in Magma shows that $\mathcal{X}_{0}^{\prime}$ is regular and that the fibre over $q$ has only one component. Let $\mathcal{X}^{\prime}\to\mathcal{X}$ be the base change of $\mathcal{X}_{0}^{\prime}\to\mathcal{X}_{0}$ from $\mathbb{Z}$ to $\mathcal{O}_{K}$. Then the only bad fibres of $\mathcal{X}^{\prime}$ are at the two primes of $\mathcal{O}_{K}$ dividing $q$, and for $\mathfrak{q}|q$ we have $\mathcal{O}_{K,\mathfrak{q}}\cong\mathbb{Z}_{q}$ and thus $\mathcal{X}\otimes\mathcal{O}_{K,\mathfrak{q}}\cong\mathcal{X}_{0}\otimes\mathbb{Z}_{q}$. Hence $\mathcal{X}^{\prime}$ is regular and the fibres over $\mathfrak{q}|q$ still have only one component. Therefore, all $\mathcal{O}_{K}$-points have the same reduction type with respect to the model $\mathcal{X}^{\prime}$. ˜5.1 implies the existence of a nonzero log differential $\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ such that the function
$\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}},\quad P\mapsto\int_{P_{0}}^{P}\omega$
vanishes on $\mathcal{Y}(\mathcal{O}_{K})$. There is no constant in the Chabauty function since the base point $P_{0}=(0,1)$ is itself an $\mathcal{O}_{K}$-integral point. The annihilating differential $\omega$ is a nontrivial linear combination of the three basis differentials $\omega_{0},\omega_{1},\omega_{2}$ of $\mathrm{H}^{0}(X,\Omega^{1}(D))$, where $\omega_{j}=x^{j}\frac{\mathrm{d}x}{y}$. Using the known integral points $P_{0},P_{1},P_{2}$, we find that the function $\rho\colon\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}}$ defined by
$\rho(P)\coloneqq\det\begin{pmatrix}\int_{P_{0}}^{P_{1}}\omega_{0}&\int_{P_{0}}^{P_{1}}\omega_{1}&\int_{P_{0}}^{P_{1}}\omega_{2}\\[5.69054pt] \int_{P_{0}}^{P_{2}}\omega_{0}&\int_{P_{0}}^{P_{2}}\omega_{1}&\int_{P_{0}}^{P_{2}}\omega_{2}\\[5.69054pt] \int_{P_{0}}^{P}\omega_{0}&\int_{P_{0}}^{P}\omega_{1}&\int_{P_{0}}^{P}\omega_{2}\end{pmatrix}$ (7.5)
vanishes on $\mathcal{Y}(\mathcal{O}_{K})$. We remark that it does not necessarily vanish on $K$-rational points. For example for $P=(33/20,4073/8000)$, which is not integral at $2$ and $5$, one has $\rho(P)\neq 0$ in $\mathbb{Q}_{7}$. Expanding the determinant in (7.5) along the last row, we see that an annihilating log differential is given by $\omega=\alpha_{0}\omega_{0}+\alpha_{1}\omega_{1}+\alpha_{2}\omega_{2}$ with coefficients
$\alpha_{0}$ $=6\cdot 7^{2}+6\cdot 7^{3}+3\cdot 7^{4}+7^{5}+5\cdot 7^{6}+4\cdot 7^{7}+O(7^{8}),$
$\alpha_{1}$ $=3\cdot 7^{4}+3\cdot 7^{5}+5\cdot 7^{6}+5\cdot 7^{7}+O(7^{8}),$
$\alpha_{2}$ $=6\cdot 7^{2}+5\cdot 7^{3}+3\cdot 7^{4}+6\cdot 7^{5}+3\cdot 7^{7}+O(7^{8}).$
We can analyse the zeros of $\rho$ on each residue disc by computing the power series expansion in a uniformising parameter. For example, consider the point $P_{0}=(0,1)$. Its residue disc is parametrised by $Q(t)=(x(t),y(t))$ with $x(t)=7t$, $t\in\mathbb{Z}_{7}$, and
$y(t)=\sqrt{f(x(t))}=\sum_{n=0}^{\infty}\binom{1/2}{n}(f(7t)-1)^{n}=1+\frac{1}{2}(f(7t)-1)-\frac{1}{4}(f(7t)-1)^{2}+\ldots$
In terms of the parameter $t$, the function $\rho(Q(t))=\sum_{j=0}^{2}\alpha_{j}\int_{P_{0}}^{Q(t)}x(t)^{j}\frac{\mathrm{d}x(t)}{y(t)}$ is given by
$(7^{3}+3\cdot 7^{5}+5\cdot 7^{6}+O(7^{7}))t+(6\cdot 7^{4}+3\cdot 7^{5}+2\cdot 7^{6}+O(7^{7}))t^{2}+(7^{5}+2\cdot 7^{6}+O(7^{7}))t^{3}+\ldots$
The sequence of valuations of the coefficients starts $(\infty,3,4,5,6,11,8,\ldots)$. Using the fact that the basis differentials $\omega_{i}$ have power series expansions in $7\mathbb{Z}_{7}[\![7t]\!]\mathrm{d}t$, one can bound the subsequent valuations by $8$ from below. Hence, by Strassmann’s Theorem, the only zero of $\rho(Q(t))$ in $\mathbb{Z}_{7}$ is at $t=0$. This shows that $P_{0}=(0,1)$ is the only $\mathcal{O}_{K}$-point of $\mathcal{Y}$ in its residue disc. Each of the remaining residue discs contains exactly one of the known $\mathcal{O}_{K}$-points, and the same argument shows that in each case, there are no other $\mathcal{O}_{K}$-points in that disc, proving ˜7.5.
The Sage code for this example can be found at https://github.com/martinluedtke/AffChab1.